### v1

# A Large-Scale Privacy Assessment of Android Third-Party SDKs

Abstract.
Third-party Software Development Kits (SDKs) are widely adopted in Android app development, to effortlessly accelerate development pipelines and enhance app functionality. However, this convenience raises substantial concerns about unauthorized access to users’ privacy-sensitive information, which could be further abused for illegitimate purposes like user tracking or monetization. Our study offers a targeted analysis of user privacy protection among Android third-party SDKs, filling a critical gap in the Android software supply chain. It focuses on two aspects of their privacy practices, including data exfiltration and behavior-policy compliance (or privacy compliance), utilizing techniques of taint analysis and large language models. It covers 158 widely-used SDKs from two key SDK release platforms, the official one and a large alternative one. From them, we identified 338 instances of privacy data exfiltration. On privacy compliance, our study reveals that more than 30% of the examined SDKs fail to provide a privacy policy to disclose their data handling practices. Among those that provide privacy policies, 37% of them over-collect user data, and 88% falsely claim access to sensitive data. We revisit the latest versions of the SDKs after 12 months. Our analysis demonstrates a persistent lack of improvement in these concerning trends. Based on our findings, we propose three actionable recommendations to mitigate the privacy leakage risks and enhance privacy protection for Android users. Our research not only serves as an urgent call for industry attention but also provides crucial insights for future regulatory interventions.
1. Introduction
The advent and ubiquity of third-party Software Development Kits (SDKs) in Android app development herald a profound shift in the mobile software development paradigm, transforming efficiency and accelerating the pace of innovation. These SDKs provide pre-built reusable code that offers a broad spectrum of functionalities, from fundamental features like user authentication (authe_lib, ) and data encryption (encryption_lib, ) to specialized components like analytics (analytics_lib, ), advertisement delivery (advertisement_lib, ), and user interface enhancements (UI_lib, ). The open-source nature and global reach of the Android ecosystem further fuel the proliferation of these SDKs.
Integrating an SDK into an app is often as simple as declaring dependencies in the build configuration file, thus drastically reducing development time, and allowing developers to focus on app-specific features. Nowadays, the majority of Android apps incorporate multiple third-party SDKs, with an average app integrating approximately 8.6 SDKs (salza2018, ; zhang2018detecting, ). However, this widespread dependency also introduces significant security and privacy concerns. The complexity of the software supply chain and the opaque data handling practices of some SDKs can lead to privacy risks and security vulnerabilities, either due to malicious intent or unintentional oversights. These can be exploited by malevolent actors or lead to inadvertent privacy violations, as evidenced by recent high-profile incidents related to third-party SDKs (priv_leak_blackkite, ; priv_leak_mahajan, ).
The privacy issues caused by SDKs could be further magnified in the pre-installed apps on Android devices. These apps are bundled by device manufacturers or network carriers, and therefore, often have an extremely great user population. They come with high privileges and are typically non-removable, such that an SDK embedded by them can stealthily perform higher-privileged operations that normal apps are unable to. An illustrative example is an app named “Mobile Services Manager” pre-installed on certain Android devices, which was found to download apps without user consent due to a malicious third-party SDK embedded (htfma2023msm, ; mollah2022mobile, ).
Assessing privacy practices of SDKs includes two key aspects, i.e., data exfiltration and behavior-policy compliance (or privacy compliance). The data exfiltration involves the behaviors of locating sensitive data in publicly accessible places, such as the publicly accessible directories in file systems and the Internet. The opacity of SDKs’ operations and their “black-box” nature hinders meaningful assessments, leaving both end-users and app developers in the dark and elevating risks. Dishonest or malicious SDKs may disguise as their embedding apps to extensively collect privacy-sensitive data, while the end-users are unable to distinguish which party requests their privacy data. Incidents like the X-Mode controversy (keegan2022sold, ), where user locations were tracked by over 100 apps through embedded SDKs, underscore this point and serve as a stark reminder of the challenges associated with preserving user’s privacy in third-party SDKs.
On the regulatory front, data protection regulations have been put in place globally, such as the General Data Protection Regulation (GDPR) (gdpr2016, ) and the California Consumer Privacy Act (CCPA) (ccpa2023california, ), which impose a stringent requirement on data processors. In response to this, software developers are required to provide a privacy policy to explicitly disclose what personal information is collected and how these data are handled by their software. Dissecting whether the data handling behaviors of SDKs comply with the claims in their privacy policies, which we define as behavior-policy compliance (or privacy compliance), becomes an urgent task.
Our Work
In this work, we conduct a large-scale study on the privacy practices of existing third-party SDKs. Our study covers 158 prevalently-installed SDKs from two mainstream release platforms, including Google Play SDK Index (google_play_sdk_index, ) and CAICT SDK Index (caict, ). The former is the official source of Android SDKs as a part of Google Play infrastructure, which lists the most widely used commercial SDKs in the Android app ecosystem (rodriguez2024sharing, ). The latter is claimed to be the largest release platform of mobile SDKs in China, serving as a complementary source of SDKs due to the absence of Google’s service in the Chinese market. Based on the statistical data provided by the two platforms, over 100 collected SDKs have been embedded by at least 100 apps with over 10 million installations, demonstrating the broad impact of this study.
Our study comprises two main strategies, i.e., conducting a data exfiltration assessment through static program analysis, and assessing privacy compliance through interpreting privacy policies. Specifically, we conduct a static taint analysis for tracking the flow of privacy-sensitive data within the SDK. We explore the traces that read and share privacy-sensitive data in public spaces, and accordingly investigate whether these traces can lead to potential privacy-sensitive data exfiltration. For the privacy compliance assessment, we employ a large language model (LLM) to carry out a text analysis of the SDKs’ privacy policies to identify what data is requested, cross-check the data claimed in privacy policies with the data access behaviors caught during the taint analysis, and determine if an excessive collection or over-claiming issue exists.
Our study reveals critical privacy shortcomings in Android third-party SDKs. Astonishingly, 346 traces are found to read privacy data and share them in publicly accessible places. Among these traces, 338 are confirmed to be subject to data leakage. Only 109 out of the 158 examined SDKs provide privacy policies, among which approximately 37% over-collect private data and over 88% falsely claim their data collection scopes, seriously violating global data protection norms. Our re-inspection after 12 months shows no meaningful improvement in these practices over time. We also uncover a novel system settings injection issue: some malicious SDKs exploit pre-installed apps to share privacy-sensitive data into Android OS’s public system settings, effectively creating a unique user-unresettable identifier (UUI) accessible to any app. This unprecedented finding highlights an urgent need for improvement in SDK development practices, permission management, and privacy violation detection. To this end, we further propose three key strategic recommendations as mitigation.
Contributions
Our work makes the following key contributions:
  • Detecting Potential Privacy Data Exfiltration on a Large Scale. We collect 158 SDKs from two mainstream repositories and leverage taint analysis to track sensitive data flow, thereby exploring their potential data exfiltration. We managed to find 338 data exfiltration evidenced by sharing obtained privacy data into public spaces. Our work highlights the substantial privacy risks stemming from the widespread use of third-party SDKs in the Android ecosystem, unveiling complex and often opaque data-handling practices.
  • Revelation of Privacy Compliance Landscape at the SDK level. Our study uncovers a startling level of data over-collection among third-party SDKs, with an estimated 37% of SDKs engaging in this practice. Through our re-inspection, we depict the trends of SDKs’ practices in collecting and sharing privacy data, and underscore the urgency of tackling privacy risks in the Android ecosystem at a fine-grained SDK level.
  • Discovery of Privacy Risk caused by Pre-installed Privileges. We find a system settings injection issue in the pre-installed apps, in which a malicious SDK can excessively read and share privacy data into the system settings. This enables an arbitrary pre-installed app with the malicious SDK to create uniquely unresettable identifiers and share them, violating the relevant regulations about privacy protection. It is the first time such issues have been revealed and systematically investigated.
2. Background and Problem Definitions
2.1. Android SDKs and Data Collection Practices
Android third-party SDKs represent a crucial component within the broader ecosystem of Android app development. These software SDKs offer pre-written code to app developers, enabling the efficient integration of a myriad of features and functions within their applications, ranging from user interface elements to complex data analytics capabilities. By leveraging these SDKs, developers can expedite the application development process, avoiding the necessity to “reinvent the wheel” for common functionalities. Consequently, these SDKs have become an integral part of the Android development framework, enhancing productivity and fostering innovation through code reuse.
Data collection practices associated with Android third-party SDKs, however, have been a subject of scrutiny within academic and industry communities. It has been observed that some third-party SDKs, while providing utility functions, also engage in extensive data collection activities without explicit acknowledgment or consent received from the end-users. These practices tend to infringe upon user privacy at times, collecting sensitive information such as device identifiers, location data, and personal preferences. Furthermore, there are instances where data is transmitted to remote servers, raising concerns about data security and potential misuse. In light of this, there is an increasing emphasis on implementing robust privacy and security measures, such as privacy-preserving APIs and data anonymization techniques, within these SDKs to ensure ethical and lawful data collection practices.
2.2. Objectives and Problem Definitions
This work aims to (1) detect potential exfiltration of users’ privacy data existing in the third-party SDKs, (2) explore the privacy compliance of developers in the SDK release, and thereby, (3) unveil the landscape of privacy protection in the Android ecosystem at the SDK level. To this end, we propose a compliance model to systematically assess the privacy practice of Android SDK regarding personal information collection. We let $\mathbb{D}$ represent a set of diverse data considered as user privacy, and let $\mathbb{S}$ be the set of publicly released third-party SDKs. Then we define what privacy data an SDK claims to collect and actually collects as follows.
Compliance Disclosure ($\mathcal{C}$)
We let $\mathcal{C}$ denote the compliance disclosure of an SDK $\mathbf{s}\in\mathbb{S}$. Specifically, it defines the set of privacy data claimed to access in the privacy policy of the SDK $\mathbf{s}$. It shall also be assumed not to leak any privacy data to the public.
Implementation-level Practice ($\mathcal{P}$)
The compliance practice of an SDK concerns two types of operations in its implementation, i.e., collecting privacy data, and sharing the collected privacy data. We let $\mathcal{R}_{s}$ denote the scope of privacy data would be read/collected by an SDK $\mathbf{s}$ based on its implementation, in which $\mathcal{R}_{s}\subseteq\mathbb{D}$. We also let $\mathcal{U}_{s}$ denote the collected privacy data involved in any uploading/sharing operations, given that $\mathcal{U}_{s}\subseteq\mathcal{R}_{s}$. Thus, we have the compliance practice of an SDK $\mathbf{s}$ as a tuple of its collecting and sharing behaviors, written as $\mathcal{P}_{s}=(\mathcal{R}_{s},\mathcal{U}_{s})$.
Next, we formalize three types of compliance issues concerned in our investigation.
Type I: Privacy Leakage
Sharing users’ privacy in public spaces poses incompliance risks especially when the collection or the sharing operations are not performed with users’ consent. Moreover, some regulations like the GDPR have explicitly prohibited any collection of UUIs, amplifying the risks of privacy sharing on the SDK side. Here we define that any pre-recognized privacy data $\mathbf{d}$ being observed in sharing operations of an SDK $\mathbf{s}$ would constitute privacy sharing risk, detailed as follows:
(1) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},d\in\mathcal{U}_{s}\Rightarrow\,\,\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$
Besides improper sharing, recent developments in privacy protection regulations also pose stricter requirements in even reading privacy data on personal devices. Developers are required to not only declare a list of privacy data to be collected but also to ensure the actual collection operations are always consistent with what has been declared. Next, we formalize two types of inconsistency issues that could happen in privacy data collection.
Type II: Excessive Collection
Excessive collection concerns the privacy practice of an SDK that attempts to read more types of privacy data than what it requests/claims in its privacy policy. Excessive collection infringes on users’ right to be informed as the SDK collects privacy data without users’ awareness and consent. We define this type of risk as follows.
(2) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\in\mathcal{R}_{s})\land(d\notin\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$
We remark that Google has not set a formal definition of “data collection” at the SDK level as what it does for third-party apps (google2024understanding, ). Thus, we refer to the guideline of GDPR Article 13 (gdpr2016, ; wolford2024guide, ) and take a conservative strategy by considering data accessed by SDKs as collected when evaluating the excessive collection, given that the data accessed by an SDK can flow into the host apps, which can then consume or send the data out.
Type III: Over-claiming
We assert an SDK that has an over-claiming issue if it is found to claim its access to more types of privacy data than what it actually reads. Over-claiming issues, although may not substantially lead to privacy data exfiltration, seriously violate the data minimization principle stipulated in Article 5(1)(c) of GDPR. It is defined as follows.
(3) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\notin\mathcal{R}_{s})\land(d\in\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvDash_{\mathcal{P}}\mathcal{C}_{s}.$
2.3. Threat Model
We consider the exfiltration of users’ privacy data caused by malicious SDKs should follow a unified routine. To achieve such exfiltration, the malicious SDK (the attacker) gains capabilities as follows:
  • The malicious SDK is released in a public platform/market and thus, can be embedded in an arbitrary app and installed on an Android device. Without causing an advantage to the attacker, that device is not rooted and is properly maintained with the latest security update installed.
  • The malicious SDK accesses users’ privacy data through the interfaces of the Android OS with permissions granted to the embedding app, i.e., all its collection of users’ privacy data must be realized on behalf of the embedded app.
  • The malicious SDK saves or shares the collected privacy data in a public space, i.e., there exists no party (including the embedding app) colluding with the malicious SDK in sharing the collected privacy data, or saving it in private storage.
3. Scope of Privacy Data
Table 1. List of 41 pre-identified privacy data types and the distribution of the 741 source methods
Data type & class Count Data type & class Count
(C1) Chip, Cellular & Peripheral Wifi 38
Call phone 41 Wifi MACs 5
Carrier info 5 Subtotal (245)
Cellbroadcast 2 (C3) Location & Sensors
External storage 18 Camera 54
ICCID 15 Location 87
IMEI 5 Misc. sensors 140
IMSI 3 Subtotal (281)
MEID 2 (C4) Media & Software Specific
Network info 6 Android ID 2
Phone status 1 App list 7
SD card serial 4 Audio record 22
Serial 4 Google Ad ID 3
SIM info 7 Media location 1
SMS 23 OAID 9
Telephone number 12 Screen record 1
Subtotal (148) Subtotal (45)
(C2) Wireless Communication (C5) Personal Data
Bluetooth 117 Account info 6
Bluetooth call 6 Browser bookmarks 1
Bluetooth ID 2 Calendar 2
Bluetooth MAC 36 Clipboard data 2
IP address 2 Contact list 9
SIP service (VOIP) 21 Contact log 2
SSID/BSSID 17 Subtotal (22)
Ultra-wideband 1 Total 741
Includes coarse location that is determined by network, and fine location that is jointly determined by network and GPS.
Figure 1. The workflow of potential data exfiltration exploration and compliance assessment at the SDK level
Sources
We resort to the Android developer documentation, including the permissions and OS release reports, to define the list of privacy data. Specifically, we browse the latest list of permissions defined in Android OS (google2024manifest, ) to retrieve the data that request explicit permissions to access, and from which we identify privacy data, such as users’ location, contacts, and calendars. These privacy data are complemented with privacy-sensitive items recognized from the privacy or security change section of each Android OS release report, for example, the clipboard and a group of user-unresettable identifiers (UUIs) are recognized in the release document of Android 10 (google2023privacy10, ) and added to the scope of this study. In addition, we also take existing literature in personal information collection (chen2014information, ; leung2016should, ; papadopoulos2017long, ; ren2018longitudinal, ; ren2016recon, ; wang2021understanding, ) and app analysis (enck2014taintdroid, ; gibler2012androidleaks, ; qiu2015apptrace, ; sun2016taintart, ) into consideration to complement the list of privacy data. We thus include the usage of diverse sensors, including motion, light, and pressure perception, into the list of privacy data because they have been widely studied in the research community. As a result, we managed to recognize 41 types of privacy data. These 41 types of privacy data constitute $\mathbb{D}$ defined in Section 2.2, formulating the scope of our privacy compliance assessment. The list of recognized privacy data is shown in Table 1.
Categorization
We then categorize the identified 41 types of privacy data by three criteria:
  • We primarily group those privacy data that share a similar functionality or purpose. This is to ease our privacy policy based compliance assessment, because many developers may tend to only declare the data to be collected roughly by functionality (e.g., device identifiers, location, etc.) rather than accessing specific data or certain APIs.
  • We then refine our categorization by grouping those data that can be programmably accessed through APIs in the same or similar classes based on the Android developer’s documentation. This is to facilitate our next step’s taint analysis as we need to collect as many involved APIs as possible to be the source methods, which we will detail in Section 4.4.
  • We also consider the API permissions of collecting those privacy data into our categorization. Due to the historic changes in Android privacy mechanism are mainly reflected in permission rules of calling sensitive APIs, grouping personal data that involve APIs mentioned in the same privacy change can facilitate our discussion and gain insights if any non-compliance is observed. A typical example is that the invocation permission of most APIs to read device identifiers including serial, IMEI, and ICCID, are escalated in Android 10. We accordingly group these aforementioned data into one category.
Based on the criterion listed above, we further categorize them into five classes, namely chip, cellular & peripheral (C1), wireless communication (C2), camera, location, and other sensors, (C3), media and software specific (C4), and miscellaneous personal data (C5). More details of the five categories can be found in the Appendix A.1.
Ontology for Semantic Relationships
Considering the request of personal data in privacy policies may not always written in a precise manner as defined in the developer’s documentation, and sometimes certain privacy data items have multiple names or colloquial phrases, we refer to relevant research (andow2020actions, ; bui2021consistency, ) and propose an ontology to capture the semantic relationship among the 41 types of privacy data, and consequently to guide our assessment.
[…]
Besides that, we manually collect alternative names or colloquial phrases for each type of identified privacy data and treat them as synonyms ($\equiv$). For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. Thus, all the synonymous terms of a privacy data type will be considered in our assessment.
4. Methodology
4.1. Approach Overview
We propose a four-phase approach and briefly illustrate it in Figure 1. Our study begins with collecting as many Android SDKs as possible from the two mainstream release platforms, shown as Phase 1. We first retrieve all SDKs available from the Google Play SDK Index (google_play_sdk_index, ), which is the official source of SDKs according to Android developer documentation. In addition, we also resort to the CAICT SDK index (caict, ) as the supplement source due to the absence of Google Play service in the Chinese market. As a result, we have collected 158 SDKs. For each SDK, we collect its binary file, the privacy policy or relevant disclosure, and meta information from the product webpage of the release platform. We detail our approach for this phase in Section 4.2.
After collecting the SDKs, we analyze the privacy policies of the collected SDKs, from which we identify data requested by the SDKs (shown as Phase 2). Due to the varying writing quality and styles of SDK documents, conventional privacy policy analysis approaches based on semantic ontologies and specific patterns or keywords cannot exhibit the desired performance. For that reason, we resort to state-of-the-art LLMs to analyze what types of privacy data are requested in natural language. Given a comprehensive list of privacy data types, our analysis model achieves 87.4% overall precision in inferring the request of a specific type of privacy data in a privacy policy. We detail our approach for this phase in Section 4.3, and evaluate our approach in Section 5.1.
Next, we explore data collection behaviors and potential data exfiltration of the collected SDK at the implementation level (shown as Phase 3). To this end, we perform an automatic static taint analysis to assess privacy data exfiltration in Android apps. Our taint analysis aims to identify the traces in the SDK implementation that constitute a complete privacy data reading and sharing process. An SDK attempts to read and share privacy data will be assessed to be prone to privacy leakage (issue type I defined in Section 2.2). We detail the design and configuration of our taint analysis in Section 4.4 and share our taint analysis results in Section 5.2. We discover the system setting injection issues from the collected SDKs and provide more details in Section 6. We also perform an additional round of evaluation and discuss it in Section 7.
In the end, we assess the status quo of privacy compliance by cross-checking the data requested in the privacy policy and the data collection behaviors of each collected SDK (Phase 4). Through this assessment, we aim to identify the excessive personal data access and permission over-claiming issues (type II and type III in Section 2.2) from the collected SDKs. We share our findings in Section 5.3. By jointly analyzing the potential data exfiltration found in SDKs’ implementation, we unveil the landscape of user privacy protection and then provide our insights and recommendations, which are detailed in Section 8.
4.2. Collecting Android SDKs
We aim to analyze data collection from developers’ disclosure and explore the data exfiltration through the lens of static taint analysis over the collected SDKs. Furthermore, we also assess privacy compliance by cross-checking the personal data requested by the developers against the actual collection behaviors found in the SDKs at the implementation level. To this end, in addition to downloading the SDK binaries (usually in AAR and JAR formats), we also collect the privacy policies and relevant metadata released by the developers, if available.
Sourcing SDKs from Multiple Repositories
We recognize the Google Play SDK Index as the main repository to collect existing Android SDKs. Moreover, we notice the uniqueness of mainland China’s Android ecosystem due to the absence of Google Play infrastructure, and therefore, we resort to CAICT as a supplement source to collect Android SDKs. We remark that, at the moment of this study being conducted, both sources serve as indexes that include SDKs meta information and download links but do not enforce a strict regulation and code auditing mechanism like Google Play App Store. Therefore, there is no guarantee that each SDK included in these two sources comes with comprehensive documents and a complete privacy policy requested by relevant regulations such as GDPR. In CAICT’s repository, each release version of an SDK is published separately. We need to filter out the historical versions of the duplicated SDKs and only download their latest release. As a result, we managed to download 158 SDKs as of October 2022, among which 139 are from the Google Play SDK Index and 19 are released on the CAICT website. These 158 constitute our previously defined $\mathbb{S}$ (Section 2.2) for the later assessment. We also revisit the latest versions of the same group of SDKs in October 2023 and conduct an additional assessment, which will be detailed in Section 7.
Crawling Meta Information and Privacy Policies
Next, we implement a crawler with Selenium (selenium, ) to retrieve meta information and privacy policy links from the two sources. The meta information includes the SDK name, provider (developer’s identity), Maven IDs, and functional categories. Such information will be jointly used to uniquely identify SDKs during our subsequent analysis. We then collect the privacy policies from the retrieved links. We will use them to extract the list of personal data that the SDK requests.
The retrieval of privacy policies is not as straightforward as what we do for meta information because there does not exist a compulsory standard for developers to draft their privacy policies. Thus, for the SDKs from the Google Play SDK Index, we resort to the link indexed by data safety section guidance, if provided by developers, and download the linked webpage. Similarly, we download the webpage from the URL link indexed by SDK privacy policy link [^1] We resort to Google Webpage Translate to process all non-English text from the involved webpages. in the SDK details webpage for those SDKs collected from the CAICT website. Due to the low availability of privacy policies, we only managed to collect the privacy policies of 52 SDKs directly from the two platforms. For the remaining SDKs that do not provide a URL link to their privacy policies, we manually search the relevant documents on the Internet with a few keywords including “disclaimer”, “policy”, “disclosure”, and “guidance”. If the documents have been successfully found, we replenish the webpage into our privacy policy dataset. As a result, we managed to acquire valid privacy policy webpages for 109 SDKs in total. The text of privacy policies is extracted for our next step analysis.
4.3. Identifying Data Requested in Privacy Policies
After collecting the 109 privacy policies, our next step aims to identify the data requested by each SDK from them. In contrast to traditional app privacy policies that are written in user-oriented language, SDK privacy policies tend to be developer-oriented and may contain various terminology that are domain-specific (e.g., ICCID, OAID, BSSID) or technically synonymous (e.g., Google Ad ID, which also appears as Advertising ID, or GAID). This characteristic makes it challenging for existing methods that build privacy policy ontologies using trained models (e.g., Name Entity Recognition) (andow2019policylint, ; andow2020actions, ; bui2021consistency, ) and analyzing privacy policy frameworks based on specific sentence patterns (xie2022scrutinizing, ; yu2016can, ; zimmeck2016automated, ) to extract domain-specific entities accurately. Therefore, we leverage a LLM to address the technical challenges posed by existing approaches, and we resort to Natural Language Inference (NLI) techniques to infer the data collection from the privacy policy documents.
Figure 2. The workflow of our privacy policy analysis to find out what an SDK claims to collect
Adoption of NLI Model
NLI has shown promising capability in understanding the context of natural language in literature of other domains (harkous2022hark, ). In this work, we leverage the roberta-NLI model, [^2] The full name of the model is roberta-large-snli_mnli_fever_anli _R1_R2_R3-nli. We shortly write as roberta-NLI to save the space. a state-of-the-art transformer-based LLM specialized in NLI tasks (nie2020adversarial, ), to help us identify data requested from the collected privacy policies. We briefly illustrate our approach in Figure 2.
In the usage of the roberta-NLI model, an NLI task aims to determine whether one text $\mathcal{A}$, also known as premise, could imply or infer another text $\mathcal{B}$, also known as hypothesis. By providing both texts to the NLI model, it outputs predictions of three types of relationships, as detailed below:
  • Entailment: $\mathcal{A}$ can logically imply $\mathcal{B}$.
  • Contradiction: there is a clear logical contradiction between $\mathcal{A}$ and $\mathcal{B}$.
  • Neutral: $\mathcal{A}$ and $\mathcal{B}$ do not exhibit a clear entailment or contradiction relationship.
Inference Setups
We first manually formulate hypotheses based on the 41 pre-identified data types (refer to Section 3). After that, we parse all sentences from the SDK privacy policies, employ them as the premises, and pass them to the NLI model to determine the entailment relationship with the hypotheses.
The formulation of hypotheses is shown as a non-trivial process due to two challenges found during our study. The first challenge is the developer-oriented writing style. For that reason, given a data type, we have to provide as many detailed items that fall into the same scope as possible to form the object of a hypothesis sentence. For example, both the user name and email are compositions of the account information of an Android device. Moreover, we manually collect all possible technically synonymous terms for each data type to pursue satisfactory completeness of our hypotheses. For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. In that case, we formulate hypotheses for each synonymous term. Besides that, we also tailor the verbal phrase for different objects in forming the hypotheses. For each type of privacy data (including its synonyms), we select all applicable verbal phases from a predefined set of verbs including “collect”, “read”, “request”, etc.
Our second challenge is the complexity of the context in SDK privacy policies. Due to the adopted NLI model is not domain-specifically trained, it is hard to know whether a sentence from the privacy policy (premise) describes the SDK’s behaviors. For example, during our preliminary study, we found a common sentence format from many SDKs’ privacy policies, e.g., “According to GDPR, all collection of device identifiers must be declared in advance”. Conventional inference task tends to determine that sentence implies the collection of device identifiers with a high confidence score. However, it is irrelevant to the actual behaviors of the analyzed SDK. To avoid the disturbance of irrelevant context, we further tailor our hypotheses for each data type. Instead of a single hypothesis to determine if the premise mentions the collection of certain data, we devise three hypotheses corresponding to three different sentiments to regulate the contextual scope, namely positive hypothesis (shown as PH in Figure 2, e.g., “It will collect …”), negative hypothesis (NH, e.g., “It does not collect …”), and irrelevant hypothesis (IR, e.g., “It does not mention whether to collect …”). As a result, we have composed a list of 507 hypotheses covering the 41 pre-identified data types. The full list of hypotheses is provided in Appendix A.2.
[…]
Inference of Data Requested
We affirm a data type is requested by an SDK if the NLI model finds at least one sentence from the SDK’s privacy policy that satisfies the criteria stating that the involved SDK will collect that data. Since we formulate three hypotheses for the collection behavior of each combination of verbal phrase and object, we only consider a data type claimed to be collected by an SDK if there exists at least one combination of the associated object (e.g., location) and applicable verbal phrase (e.g., request) satisfying the logical condition as follows:
$(\text{Score}_{p}>T)\wedge(\text{Score}_{p}>\text{Score}_{n})\wedge(\text{Score}_{p}>\text{Score}_{i})$
where $\text{Score}_{p}$ indicates the entailment confidence score for the positive hypothesis. An entailment inference requires $\text{Score}_{p}$ to be greater than a threshold value, written as $T$. Simultaneously, $\text{Score}_{p}$ must be greater than the scores for negative and irrelevant hypotheses (i.e., $\text{Score}_{n}$ and $\text{Score}_{i}$), indicating that the premise does not only “mention” the collection of that data type but also confirm that collection behavior is “performed” by the SDK itself.
Finding the Optimal Threshold
Intuitively, to infer a specific type of data being collected, the entailment prediction score ($\text{Score}_{p}$) is supposed to be greater than 0.5 to ensure it is always higher than the prediction of the neutral and contradiction hypotheses. However, our preliminary experiments observe a large number of false positives by setting the threshold to 0.5. For that reason, we perform a small-scale analysis to tune the threshold value from 0.5 and above until we find an optimal value that minimizes the number of false positive predictions. Specifically, we only consider SDKs developed by Google from the Google Play SDK Index based on our assumption that those SDKs should provide high-quality privacy policies. We gathered 16 SDKs authored by Google, among which 13 SDKs provide privacy policies. Due to shared privacy policies among some SDKs (i.e., SDKs for Google Firebase), we eventually reviewed five distinct privacy policies. We manually analyzed declared privacy data types from the five privacy policies to set a ground truth. We then fine-tuned the threshold value and discovered that the NLI model could achieve perfect prediction with a threshold of 0.73. For that reason, we adopt 0.73 as the optimal threshold in our large-scale privacy policy analysis.
To validate NLI model performance on analyzing privacy policies of all the collected SDKs, three members from our organizations are invited to perform a manual validation independently. Each of them holds expertise in Android app analysis and development, with one member additionally having proficiency in drafting privacy policy documents. We will discuss the evaluation outcomes of our inference in Section 5.1.
4.4. Exploring Privacy Collection and Potential Leakages
We leverage the static taint analysis techniques to explore privacy collection behaviors and potential exfiltration in SDKs. We assume that a valid privacy data exfiltration is constituted of a collection action that queries one or more privacy data from the OS, and a sharing action that shares the obtained privacy data, directly or indirectly, to the public spaces including the Internet, system settings, and the public storage. Thus, we stipulate an SDK conducts privacy collection if there exists a data flow that starts from a privacy data collection action in its implementation, and this privacy collection would result in a data exfiltration if the data flow sinks at a sensitive sharing action.
Configuring Taint Analysis
After confirming the list of privacy data, we retrieve the API documents to collect all relevant APIs that contribute to a collection action and include them into the source methods for taint analysis. We refer to an existing research (rasthofer2014machine, ) which proposes a technique named SuSi, which manages to identify and classify over 900 API methods that can potentially leak users’ privacy in Android 4.3 and earlier versions. Considering there has been a decade’s evolution of Android OS since the release of SuSi, many APIs are no longer valid because they have become deprecated and then removed from later OS releases. Meanwhile, new hardware and functionalities, e.g., motion sensors and ultra-wideband (UWB), have emerged during the rapid evolution of the Android ecosystem, and accordingly, the scope of privacy data keeps being extended. For those reasons, we first map the APIs identified in (rasthofer2014machine, ) with the 41 types of privacy data. We iterate the developers’ documents of those APIs and filter out those APIs that are no longer valid in Android Open Source Project (AOSP) implementation. We then manually visit the developer’s document and AOSP source code to collect as many APIs that are associated with the new data types as possible. As a result, we managed to collect 741 source methods. [^3] We count the number of methods by unique method signatures in this paper. The distribution of 741 methods corresponding to the 41 types of privacy data can be found in Table 1.
Table 2. A brief statistics of recognized sink methods
Purpose Class name Count
Network Transfer java.net.HttpURLConnection 1
java.net.URL 2
java.net.URLConnection 2
org.android.spdy.SpdyRequest 12
org.android.spdy.SpdySession 3
org.apache.http.client.methods.HttpGet 8
org.apache.http.client.methods.HttpPost 8
org.apache.http.params.HttpParams 1
File Saving java.io.Writer 5
java.io.FileOutputStream 1
System Settings android.provider.Settings 3
Total 46
Next, we retrieve all methods that perform a sharing action from Android SDKs. Although there are not many new APIs introduced after Android 4.2 that can be treated as sink methods, we find most sink methods studied in earlier literature including SuSi (rasthofer2014machine, ) are no longer valid in the latest Android versions and therefore be excluded in this study. For example, logging is no longer a feasible approach for sharing data among different apps because an app cannot read logs created by another (google2023loginfo, ). Similarly, methods to access and manage SQLite databases are not counted because the database files are protected by the sandbox mechanism of Android and are not visible to other apps (google2022sandbox, ). In the end, we identify 46 methods from 11 classes and then configure them as sink methods. We present brief statistics of the sink method by their containing classes in Table 2.
The majority of sink methods are relevant to data transfer via network interfaces, reserving 37 out of 46 identified method signatures. Besides that, five methods defined in java.io.Writer and one method from java.io.FileOutputStream are recognized as sink methods that can share sensitive data via public storage. The remaining three sink methods are counted because they can be used for modifying the system settings, which is known as an undocumented channel to share data among different parties on the device (meng2023post, ). We note that writing contents into the system settings requires privileged permission and we will discuss it in Section 6.
Adopting Automatic Taint Analysis
We learned that there are some off-the-shelf automatic static taint analysis tools for Android apps. Our preliminary investigation considers three open-sourced tools, namely FlowDroid (arzt2014flowdroid, ), Mariana Trench (mariana_trench, ), and AppShark (appshark, ). We set up these three tools on our experimental PC, feed four best-selling non-game apps from the Google Play App Store [^4] The four apps used for our preliminary investigation are Facebook, Instagram, Snapchat, and TikTok. into these tools, and explore data flows simultaneously containing a privacy collection action and a sharing action. Our observation shows that AppShark performs the best in terms of time efficiency, success rate of analysis, and capability of handling inter-process communication (IPC). For those reasons, we adopt AppShark for the taint analysis of SDKs on a large scale.
5. Evaluation
Our evaluation aims to assess the performance of our proposed NLI model-based privacy policy analysis and explore the privacy data collection and potential leakages in the collected SDKs. After that, we analyze the landscape of privacy preservation at the SDK level by cross-checking the data requested in the privacy policies and the data collected in SDKs’ implementation.
Implementation
We implement our privacy policy analysis in Python with PyTorch. The NLI model is provided by a Python package named Transformers. [^5] Can be accessed through URL https://huggingface.co/docs/transformers/index Our static taint analysis is implemented in Java and it supports users to specify target apps and customize rules for taint analysis in a configuration file. For each SDK, we implement a blank Android app to embed it and pass the app package to the automatic taint analysis program. Considering that taint analysis may unavoidably produce false positives, manual validation of the tainted traces is performed to ensure the correctness of our further assessment. We will open-source our implementation and relevant artifacts upon paper acceptance.
Evaluation Goals
Our evaluation aims to answer the following three research questions (RQs):
  • RQ1. Can our privacy policy analysis identify the privacy data claimed to be collected by SDKs on a large scale? What is the performance of our NLI model-based approach?
  • RQ2. Can our taint analysis find privacy data collection and sharing behaviors from the SDKs? How many tainted traces are detected? Can those tainted traces (really) leak users’ privacy?
  • RQ3. By cross-checking their privacy policies, what is the status quo of privacy data protection at the SDK level? Do the collected SDKs comply with their privacy policies?
5.1. RQ1: Privacy Policy Analysis
Table 3. Evaluation outcomes of privacy policy analysis (only data types requested are displayed)
* We use a broad sense of device identifiers to represent all hardware UUIs, including IMEI, MEID, IMSI, ICCID, and serial.
We perform NLI model-based analysis on the 109 SDK privacy policies and present the analysis outcomes in Table 3. Our analysis manages to identify 214 data requests from privacy policies. Among the 214 data requests, IP address reserves the largest share, contributing to 47 (21.9%) occurrences. network information and location are the remaining data types among the top three, which are requested in 30 (14.0%) and 28 (13.0%) occurrences, respectively. From the perspective of SDKs, the number of data types claimed by an SDK ranges from 0 to 6, with the average number 1.9.
We also conduct a manual validation to evaluate the performance of our proposed NLI model-based analysis in identifying data requested from privacy policies. After an independent manual inspection of our findings in the 109 privacy policies, we validate 187 data requests covering 19 data types, and accordingly treat these 187 requests as the ground truth in evaluating our analysis approach. Overall, the evaluation outcomes show that our approach is capable of performing privacy policy analysis on a large scale, evidenced by a promising precision of 87.4% and F1 score of 89.5% (see Table 3).
We find our NLI model-based analysis achieves perfect precision in detecting data collection for 10 data types despite 27 false positive cases being observed, i.e., wrongly recognized by the NLI model. Our evaluation outcomes show that account information (6), network information (5), and SMS (4) are the top three data types involved in false positives, indicating our NLI model together with the formulated hypotheses cannot well infer the collection of certain data types. The main reason for these false positives may be the misunderstanding of the adopted LLM to the terms we use to describe the involved privacy data, as well as the scope of certain data types. A typical example is the analysis of collecting privacy data relevant to SMS. Our approach is capable of understanding the literal meaning of SMS but fails to differentiate SMS and other messaging approaches such as e-mails. We find that our adopted NLI model mistakenly entails the collection behavior with high confidence when processing sentences like “we may send email to the users…” in some SDKs. We elaborate more on these false positives with real examples in Appendix A.3.
5.2. RQ2: Privacy Collection and Potential Leakages at the SDK level
Our static taint analysis manages to identify 2318 occurrences of privacy data access in their implementation, covering 27 out of the 41 pre-identified data types. 95 out of 158 (60.1%) SDKs are found to read privacy data, and 34 (21.5%) SDKs have at least one trace tainted. Among these access behaviors, 346 (14.9%) accesses are tainted during the analysis, indicating their data flows contain a sharing action that potentially leaks the accessed privacy data to the public spaces. We present the taint analysis outcomes by data types in Table 4. We also performed manual validation to exclude false positive tainted traces, which we will elaborate on with examples in the remainder of this section.
Table 4. A brief statistics of the taint analysis results (only data types called and/or tainted are displayed)
Distribution of Data Collected & Shared
As shown in Table 4, location and sensors’ data (category C3) reserves the greatest portion of collection actions (reflected in the total number of caught access traces in the Table), contributing 1002 out of 2318 (43.2%) caught collection actions. The remaining privacy data categories ordered by their number of caught collection actions are data for wireless communication (C2, 22.8%), data for chip, cellular and peripheral (C1, 17.9%), media and software-specific data (C4, 13.4%), and personal data (C5, 2.8%), respectively.
From the privacy data sharing perspective, the access of wireless communication data (C2) constitutes the largest part of the tainted flow, as 172 out of 346 (49.7%) traces are tainted. The numbers of sharing actions involving data in categories C3 and C1 are following up, reserving 37.6% and 12.7% of all sharing actions, respectively. No sharing actions about data in categories C4 and C5 are observed. From the identifiability perspective, we find that most privacy data involved in sharing actions are referred to as personally identifiable information (PII) in relevant literature (ren2018longitudinal, ; ren2016recon, ; shen2021understanding, ; wang2020exploring, ).
We remark that whether SDKs can collect privacy data eventually depends on the permission and privilege of their embedding apps, especially for the collection and sharing of PIIs. We also highlight that the tainted traces do not automatically imply the potential exfiltration in runtime, which we will discuss later in this section.
1public static String a(Context context) {
2 ConnectivityManager cm = (ConnectivityManager) context.getSystemService(Context.CONNECTIVITY_SERVICE);
3 NetworkInfo info = cm.getActiveNetworkInfo();
4 if (info.isConnected()) { return "wifi"; }
5 ...
6 try {
7 return info == null? "" : info.getTypeName()+"-"+info.getSubtypeName()+"-"+info.getExtraInfo().toLowerCase();
8 }
9 catch(Exception e) { return ""; }
10}
Figure 3. A false positive example of SSID reading (line 7)
Potential Privacy Data Leakage (Type I Issues)
During our analysis, we noticed that static taint analysis cannot well determine the execution paths in conditional and branched code. If any tainted trace is found spanning over an unreachable, dead, or logically infeasible code, we exclude the tainted trace from the statistics of potential leakages. In Figure 3, we provide an example code reverse-engineered from one of our examined SDKs named “mipush”. We find it attempts to read the SSID of the connected WiFi but is later assessed by us as a false alarm. The SSID reading action (can be realized by the method getExtraInfo (line 7) until Android 10) would only take place if the device is not connecting to WiFi (after line 4), therefore we determine this is a logically infeasible code regarding the privacy data access. In this study, we manually exclude the false alarms from the 346 traces and manage to identify 338 (97.7%) logically valid traces, among which each trace indicates an occurrence of privacy leakage issue.
For the sharing operations, we observe that 332 out of 338 (98.2%) valid traces are tainted at network interfaces. The six remaining traces are tainted at system setting APIs. Compared with sharing privacy data through the Internet, we remark that writing privacy data in system settings is a more worrying issue because it exploits the privilege of pre-installed apps and is capable of sharing user privacy even without reading any privacy data from the Android OS. We will detail this in Section 6.
5.3. RQ3: SDK-level Privacy Compliance
We resort to cross-checking the privacy policies with the actual data collection behaviors of the collected SDKs to understand the status quo of privacy compliance at the SDK level. We remark our assessment only focuses on the 109 out of 158 (69.0%) SDKs that provide privacy policies, and the 27 data types that appear in the data collection behaviors observed from the taint analysis. We present our assessment results in Table 5.
Table 5. Assessment of privacy compliance by data types
Excessive Collection (Type II Issues)
Our evaluation finds that 40 out of 109 (36.7%) SDKs collected excessive privacy data compared with their privacy policies (e.g., an SDK collects personal identifiers but does not declare them in its privacy policy). Among the 40 SDKs that excessively collect users’ privacy, the number of undeclared data types ranges from 1 to 9. Each SDK has an average of three undeclared data types. As shown in Table 5, 22 out of the 41 pre-identified data types (81.5%) have been excessively collected by the examined SDKs. The app list, clipboard data, and location are the top three most excessively collected privacy data types. Specifically, 22 SDKs (20.2%) collect the app list without any declaration in their privacy policies, followed by clipboard data and location that are excessively collected by 11 (10.0%) SDKs. By jointly considering data collection behaviors (refer to Table 4), our findings show that the app list and clipboard data are the two most stealthily collected data types. Only 4 SDKs collecting the app list have mentioned it in their privacy policies (leading to 22 excessive collection behaviors). None of the 11 caught SDKs have declared to read the clipboard data.
Over-claiming (Type III Issues)
Apart from the excessive collection, we find that the over-claiming issue in privacy policies is far more pervasive in the tested SDKs. In our study, 96 out of 109 (88.1%) SDKs are found to claim more data types in privacy policies than what they (really) need. The IP address, network information, and MEID are the top three data types that have been over-claimed in the tested SDK privacy policies. Especially, 44 out of 109 (40.4%) SDKs that provide privacy policies over-claim their access to IP addresses. Such pervasive claims, again, reflect the worrisome landscape of privacy compliance at the SDK level.
Lessons Learned
In addition, our study unveils two interesting facts. First, we find that some SDK developers tend to use vague or general terminologies to describe collected privacy data such as device identifiers, advertising identifiers, and device information instead of listing out specific data collected. Developers’ interpretation of these ambiguous terminologies greatly differs. For example, a few developers define “identifiers” as advertising identifiers although the majority use this terminology to represent device identifiers, resulting in over-claiming cases in device identifiers but excessive-claiming cases in Android ID and Google Ad ID. The second fact is that many developers fail to clearly specify the sources of collected data types, especially the location. Commonly, location can be obtained by accessing GPS in mobile devices or inferred from the information of the connected network and IP addresses. Our analysis of privacy policies finds that many SDKs tend to declare their access to both the GPS and IP addresses to determine locations. However, our taint analysis shows that most SDKs turn to invoke location APIs only, leading to many over-claiming cases of IP addresses. This phenomenon may confuse the end users in understanding which exact data types are collected from them, resulting in poor transparency of the privacy policies.
Responsible Disclosure
We have reported all our findings to the relevant parties and we also have kept them confidential for at least 90 days. For each SDK that found a compliance issue, we manually searched the contact left by its developers and e-mailed our findings since our first round of SDK investigation. Although we observed slight mitigation from re-inspection, we did not receive positive responses regarding non-compliance concerns from the developers that we have attempted to contact. We also expressed our concern to Google about the system settings injection issues (to be detailed in Section 6) and suggested removing the pre-installed privileges in writing system settings in the future release. However, we are told that the pre-installed privilege is not regarded as a vulnerability and the relevant permission mechanism would remain unchanged. It is at the device manufacturer’s discretion to select and grant pre-installed privileges to third-party apps.
6. System Settings Injection Issues
Figure 4. Information flow demonstrates how the device ID is accessed and saved to the system settings by one of our analyzed SDKs (the obfuscated code is obtained from reverse engineering and has been simplified to save space)
Table 6. List of system setting entries found in our investigation that contains a UUI
SDK name Entry name Description of the value Operations to System Settings
[…]
Baidu LBS com.baidu.deviceid.v2 App’s package name and Android ID, MD5 $\ddagger$ Read only
[…]
$\dagger$ depending on the permissions granted to the embedded app. $\ddagger$ inferred description based on the reading logic in the implementation.
During our analysis, we find some traces from an SDK named “Alicloud push” are sunk at writing a value into the system settings. Specifically, two entries of system settings are involved, as shown in Table 6. Writing into the system settings requires privileged permission so that ordinary third-party apps are unable to do so. However, we find that SDKs being embedded by pre-installed apps [^6] We provide more technical details of the pre-installed apps and permission for writing into system settings in Appendix A.4. can write arbitrary values into the system settings even without user’s awareness. We name this finding a system settings injection issue, because it can be manipulated by an SDK to create and share user-unresettable identifiers (UUIs).
Although leaking privacy data through covert channels has been studied in existing literature (reardon2019fifty, ), the system settings injection issues still seriously concern us as the caught SDK subtly evades accessing highly sensitive UUIs (e.g., IMEI, as discussed in  (reardon2019fifty, )) but attempts to generate UUIs on their own. The caught SDK then takes advantage of its popularity and abuses the “preinstalled” privilege by counting on itself to be embedded in pre-installed apps or even privileged system apps. By exploiting this issue, the system settings become a covert channel to enable privacy-sensitive data sharing to apps without privileged permission.
Figure 4 illustrates an example of system setting injection from our findings. The source code is obtained through reverse engineering. In order to obtain a UUI, it attempts to invoke a privileged API (line 5, left), query the existence of pre-written system properties (lines 6 and 10-15, left), and read the Android ID (lines 7 and 16-24, left) sequentially, as shown in the figure. Due to the first approach requests the embedded app to be a system app with “READ_PRIVILEGED_PHONE_STATE” permission, and the second approach is only viable when there exists a system app on the same device to write the two system properties in advance, we consider both them infeasible by default on nowadays Android devices. However, the third approach, i.e., reading Android ID through a system setting table named SECURE, does not request special permission. As a result, the SDK can at least obtain the Android ID to uniquely identify the device. It then performs a hash function over the obtained Android ID (lines 19-20, right) and writes the hash value into the system settings with a customized key (lines 25-27, right). This process, although it complies with the Android permission mechanism, transforms the value of resettable Android ID into a de facto UUI that can be accessed by any app that knows the customized key in the system settings.
Considering the exploitability of the pre-installed privileges, we carry out another round of investigation of all 158 SDKs to find if there exists any other potential privacy infringement by taking advantage of the system settings mechanism. As a result, we managed to find another two entries that have been read in an SDK named “Baidu LBS” (see Table 6). As their names imply, these two entries are obviously written by third-party privileged apps. We suspect the formation of the two entries should follow a similar routine, and accordingly reverse engineer the involved SDK and present our inferred descriptions of the two entries in Table 6. Unfortunately, we did not find any app or pre-installed app to write these entries into the system settings. Who exploited this system settings issue remains unknown to us.
Impact
We notice that the two studied SDKs are widely adopted in third-party apps, especially in the Chinese market. Therefore, we conduct a small-scale study to explore the impact of the system settings injection issues. We collect devices from 9 OEM manufacturers (8 manufacturers’ devices purchased in the Chinese market and 1 Google Pixel device purchased in the US market, please refer to Appendix A.5 for the full list), download the packages of pre-install apps, and investigate whether the two studied SDKs have ever been embedded in any of the pre-installed apps. We observe a worrisome fact that except for the Google Pixel devices, all the tested devices have at least one pre-installed app embedding the investigated SDKs. We assume that the choice of pre-installed apps should remain the same across different device models produced by a manufacturer. By referring to the latest market report (statcounter2024worldwide), these 8 OEM manufacturers reserve over 66% global Android smartphone market share. They also shipped over 94% new Android devices in China in the year 2023 (statcounter2024china, ), implying that almost all Android smartphones sold in the Chinese market pose a privacy leakage risk.
7. Privacy Compliance Re-inspection
Our study aims to unveil the comprehensive landscape of privacy compliance at the SDK level. Instead of conducting a one-shot analysis, we carry out a re-inspection to explore the trend of privacy collection among the same group of the tested SDKs. More specifically, on the basis of our large-scale analysis performed on 158 SDKs collected in October 2022, we re-visit the two sources and collect the same group of SDKs in their latest release version as of October 2023. Thus, we conduct a cross-version analysis of the privacy collection and sharing practices of the examined SDKs.
Table 7. Changes in the tainted traces observed from re-inspection after 12 months (only tainted data types are displayed)
Data type & class Changes in tainted traces Data type & class Changes in tainted traces
[…]
$\dagger$ We use “new” to indicate data types observed re-inspection for the first time.
Trends of Privacy Collection Behaviors
We assess the same group of SDKs in their latest release versions and summarize the tainted traces in Table 7. We managed to find 311 traces tainted from the latest version of 158 SDKs. Compared with our initial assessment, we observed 35 (10.1%) fewer traces that attempt to collect user privacy data and share them in public spaces.
The change between the outcomes of the two assessments mainly falls in the sources of tainted traces, which correspond to what types of privacy data have been collected and shared. As shown in Table 7, we find obvious growth from the sharing of data relevant to media and installed software (category C4), reserving over 1/3 of the total caught traces. Among the 105 tainted traces, reading and sharing the list of installed apps ranks the top among all sources, evidenced by 53 traces related to this type of privacy data. Although the access of data in that category has been pervasively observed in our initial assessment (310 traces that are not tainted, as shown in Table 4), our taint analysis did not catch any traces of sharing such data.
Additionally, we find that the location and SSID/BSSID contribute to the two greatest declines in the number of collection behaviors. We also observe an overall decrease in the collection of UUIs, including device identifiers (category C1) and identifiers of hardware components for wireless communication (category C2). By jointly considering the historical changes in permission requirements, we find that the access rules of most data that show a decrease in collection behaviors have undergone a significant update in Android 10, in which the third-party apps’ access to personally identifiable data has been strictly restricted (google2023privacy10, ). Specifically, third-party apps’ access to UUIs is prohibited. Therefore, the implementation to collect UUIs becomes invalid unless they are embedded in system apps. Permission to access data that may leak users’ physical location, such as GPS location, SSID, and BSSID, has also been escalated. A finer location permission becomes mandatory. Such permission escalation makes SDKs unable to stealthily access the aforementioned data unless the embedding apps explicitly request permission from users.
We also observe that the trends in SDKs’ privacy compliance show hysteresis along with the evolution of Android OS, given Android 10 was released in 2019. Such hysteresis could also be reflected in the data type demonstrating the largest growth of tainted traces, namely the app list. We consider SDK developers to treat the app list as an alternative to the UUIs and share it for user identification and profiling purposes. However, we find the SDK developers’ efforts in collecting user privacy lag behind Google’s strengthening privacy protection mechanism of Android OS, since access to the app list has been restricted since Android 11, which is just one year after the release of Android 10 (google2023privacy11, ). In other words, SDK’s privacy collection via the app list is infeasible on most devices nowadays. Nonetheless, given the number of tainted traces does not experience a significant change between the two assessments, SDK developers’ privacy compliance still seriously concerns us since their latest implementation still implies their persistent attempt to excessively and stealthily collect users’ privacy data.
8. Discussion
8.1. Causes of Privacy Mishandling
Privilege Free-riding
An SDK’s access to privacy data is determined by permissions granted to its embedded apps. Third-party apps can obtain privileged permissions if pre-installed by OEMs. A malicious SDK embedded in a pre-installed app can exploit the system settings injection issues and execute privilege free-riding attacks, creating and sharing UUIs via the system settings. As reading system settings requires no permissions, privilege free-riding significantly undermines Google’s privacy protection efforts in restricting third-party app access to UUIs since Android 10.
Insufficient Regulations
The release of Android apps faces strict regulation. Apps’ non-compliance in privacy preservation would result in removal from app stores or fines. However, SDK release lacks such fine-grained regulation and sufficient legislation to ensure compliance with personal data protection laws. The CAICT repository serves only as a publishing platform and search engine for available SDKs, without hosting privacy disclosures/policies. The Google Play SDK Index allows developers to voluntarily provide privacy policies but does not host binaries or enforce strict policies like its app store. Not all listed SDKs are “registered” in the Google Play SDK Console, meaning no guarantee of compliance with basic terms and conditions. Notably, even some Google in-house SDKs lack privacy policies on the SDK Index.
Furthermore, there is no platform-level compliance checking mechanism to ensure complete permission disclosure. The SDK Index only displays reminders about privacy-sensitive permissions, implying developers must undertake ultimate responsibility for privacy protection, regardless their code is self-implemented or imported from external parties.
8.2. Our Recommendations
Precautions of Privilege Abuse
We regard that the pre-installed privileges open an attack surface for a malicious SDK, specifically when the SDK uses app-differential Android ID to identify a device user, transforms it into a de facto UUI, and provides free access with all apps installed on the victim device. Based on this finding, we advocate Google restricting pre-installed apps from writing data into the system settings in the future. In fact, pre-installed apps are free to create a customized UUI and share it even if it has not read any privacy data. Stop automatically granting WRITE_SETTINGS and instead allow apps with “preinstalled” privilege to explicitly request it at runtime can mitigate the risk of system settings injection. We also advise OEM manufacturers to minimize the number of pre-installed apps on their devices.
Strengthened Platform Regulation
Currently, both release platforms selected in this study are merely index websites, which do not host the SDK binaries and lack substantial enforcement. Developers voluntarily register SDKs on these platforms. For that reason, we call for strengthened regulation over the usage of Android SDKs. We also advocate platforms performing static analysis over the SDK binaries to exhaustively list requested permissions, especially for those at the dangerous level and above. This measure could address the black-box concern for app developers.
Transparent App Development
We advocate transparent management over the permission requested by the embedding SDKs. Thus, app developers can have their own discretion in granting permissions without compromising functionality. We are delighted to see Google’s efforts in improving the SDK’s governance. A new mechanism called SDK Runtime has been introduced in Android 13 that provides a dedicated execution environment for SDKs (google2024sdk, ). However, this mechanism is not compulsory and still faces many usability and functionality limitations at the moment of this work being performed. Besides that, how to ensure the cooperation of SDK developers remains a challenging issue, given many SDKs are released and maintained in non-Google platforms.
8.3. Limitations
Absence of Dynamic Testing
While our static taint analysis approach is effective, dynamic testing can improve efficiency by automatically identifying and excluding logically invalid tainted traces, reducing manual validation efforts. However, dynamic SDK testing is challenging due to inadequate documentation, undisclosed interfaces, and highly event-driven/credential-required logic. We aim to explore potential dynamic testing solutions in the future.
Coverage of Targeted Methods
Our analysis covered 741 API source methods and 46 sinks. Although we frequently update the configuration during this study, we may still miss some data exfiltration channels as new APIs can be introduced along with OS evolution and emerging hardware.
Elusive and Ambiguous Disclosure
Although developers should clearly disclose personal data collection, we find that elusive or ambiguous phrasing widely exists in many SDK’s privacy policies, leading to potential false negatives in our NLI model-based analysis. We aim to explore the latest advancements in LLMs and improve our approach to minimize such impacts.
Scope of SDKs
Our assessment covered 158 Android SDKs from two mainstream release platforms. However, privacy data leakage can occur through non-Android APIs and even pure Java logic, suggesting our future work should extend beyond Android SDKs to any libraries or tools integrated into Android apps capable of handling personal data.
9. Related Works
Privacy Compliance Evaluation
Numerous works focus on detecting inconsistencies between the privacy practices of an application and its descriptions or its privacy policies (bui2023detection, ; gorla2014chabada, ; liu2018large, ; pandita2013whyper, ; qu2014autocog, ; slavin2016toward, ; wang2018guileak, ; zhou2022policycomp, ). To automate the risk assessment of Android applications, WHYPER (pandita2013whyper, ) and AutoCog (qu2014autocog, ) leverage NLP techniques to evaluate the necessity of required permissions based on applications’ descriptions. Gorla et al. (gorla2014chabada, ) proposes a prototype to identify mismatches between app descriptions and app behaviors by clustering Android apps based on their description topics and identifying outliers in each cluster with respect to their API usage.
More recently, there have been several works examining third-party data harvesting, user consent for data sharing, service misconfigurations, and privacy policy conflicts in Android apps. XFinder (wang2021understanding, ) uses dynamic analysis and NLP techniques to parse terms-of-service and extract data-sharing policies. Nguyen et al. (nguyen2021share, ) studies user consent verification through traffic analysis and ablation experiments. Zhang et al. (zhang2020does, ) highlights the issue of analytical service misconfigurations, while Yu et al. (yu2018ppchecker, ) scrutinizes the apps’ privacy policies with regard to the inclusion of third-party libraries. Andow et al. (andow2020actions, ) proposes POLICHECK that identifies ambiguous or omitted disclosures of third parties. Despite these efforts, existing tools often neglect to assess the compliance of third-party libraries themselves with regulatory requirements. In comparison, our work focuses on compliance of third-party SDKs, checking the practical end-to-end data handling against declarations of developers. We also specifically consider the elevated risk of privacy leakage in the context of pre-installed apps.
Privacy Leakage Assessment
There is a substantial body of literature addressing various aspects of privacy concerns in the Android ecosystem (enck2014taintdroid, ; zhang2009peep, ; zhou2013identity, ). Recently, Meng et al. (meng2023post, ) studied user-unresettable identifier safeguards on a wide range of Android devices. They unveil widespread privacy issues, emphasizing the necessity for robust device-level data protection regulations. He et al. (he2018dynamic, ) utilizes dynamic analysis to explore privacy leaks from third-party libraries in Android apps, focusing on permission-related personal data. Their methods achieve a high real-time detection accuracy across 150 popular apps. Ekambaranathan et al. (ekambaranathan2021money, ) focuses on children’s app data usage and disclosure, gaining insight from surveys and interviews with app developers. Liu et al. (liu2019privacy, ) examines data leakage from nine analytic libraries across 300 apps, using both static and dynamic analyses. Razaghpanah et al. (razaghpanah2018apps, ) detects third-party advertising and tracking services via dynamic analysis of network traffic data. However, it identifies limited types of personal data and hardware identifiers. Our work comprehensively investigates privacy compliance at the SDK level on a large scale, complementing existing research to unveil the privacy preservation landscape in the software supply chain of the Android ecosystem.
10. Conclusion
In this study, we have delved deep into the intricate privacy challenges associated with the widespread use of third-party SDKs in Android apps. Our large-scale analysis involved an evaluation of 158 third-party SDKs, prevalent across Android applications, with an approach intertwining rigorous taint analysis for data collection practices and LLM-aided interpretation of privacy policies. The findings were startling, with 338 potential privacy leakages detected from 158 SDKs. Besides that, our study reveals that less than 70% of examined SDKs provide privacy policies, among which approximately 37% of SDKs are found indulging in the over-collection of privacy data, signaling a clear violation of privacy norms. Based on our findings, we believe that tighter regulatory measures, improved development guidelines, and transparent permission management are essential steps toward mitigating privacy risks in Android applications.
Appendix A Appendix
A.1. The Five Categories of Privacy Data
(C1) Chip, Cellular & Peripheral
Each Android smartphone device has several unique identifiers that are associated with its hardware components including the motherboard, cellular chip, and peripheral hardware (e.g., external storage and SD card). These identifiers, such as serial, IMEI, and IMSI, are usually user-unresettable. Moreover, the invocation of some core functionalities associated with the cellular chip are also considered privacy sensitive (e.g., dialing a phone call, and sending an SMS), and therefore are categorized in this class, too. Apps’ access to these data often requests privileged permissions. Starting from Android 10, all user-unresettable device identifiers need “READ_PRIVILEGED_PHONE_STATE” permission, and therefore, are supposed to be read by system apps only (google2023privacy10, ).
(C2) Wireless Communication
. WiFi and Bluetooth are two common features of Android smartphones to perform wireless communication. Privacy data of this class include persistent identifiers such as MAC addresses, environmental volatile information that can be used to infer the user’s identity, including paired Bluetooth device ID (shown as Bluetooth ID), IP address, SSID and BSSID, and privacy-sensitive data generated via Wifi or Bluetooth functionalities, for example, making a call through Internet (SIP) or Bluetooth. In addition, since Android started third-party apps’ support of ultra-wideband (UWB) communication in version 12 and above (google2023uwb, ), we also consider data generated by UWB in this class. Considering data in this category may partially disclose users’ identity and/or location, Android has strengthened the permission mechanism for API requesting this type of data. For example, the MAC address randomization has been introduced in Android 6 and escalated restriction has been further enforced in Android 10 (google2023privacy10, ).
(C3) Camera, Location & Other Sensors
A modern Android smartphone is usually equipped with many sensors, for example, the camera, location sensor (e.g., GPS), and motion sensor. The data produced by these sensors are widely recognized as user privacy because they can be used to infer a user’s behavior, customs, and environment.
(C4) Media & Software Specific
This class concerns data associated with multimedia features and user ID generated by the OS or certain software. For example, the root directory of media files can be misused by malicious apps to save and read irrelevant files, audio recording, and screenshot functions can leak the user’s biometric voice and behavior. The software-defined ID can uniquely identify the user himself/herself but it can be reset by users. It is noteworthy that the list of apps or packages installed has been identified as privacy sensitive and can only be accessed by system apps since Android 11 (google2023privacy11, ).
(C5) Personal Data
Data falling in this class are mainly privacy-sensitive data that disclose users’ behavior or social relationships. For example, leakages of account information may disclose the social apps that the user has been logged in to. Unauthorized access to the contact list leaks users’ social relationships and can be manipulated for criminal purposes. Disclosure of clipboard data partially unveils the user’s real-time behavior, and therefore, severely infringes his/her privacy. Unlike other categories, permissions to access data in this group are highly varied. Accessing each type of data in this category usually requires unique permission.
A.2. List of Hypotheses Composed for NLI Tasks
Table 8. Combination of objects and applicable verbal phrases to form hypotheses
The complete list can be found in Table 8. Note that we compose a positive hypothesis (“It will…”), a contradiction hypothesis (“It does not…”), and an irrelevant hypothesis (“It does not mention whether to…”) for each combination of object and applicable verbal phrase. Given 169 combinations produced, we compose a total number of 507 hypotheses for the NLI tasks.
A.3. Discussion of False Positive Cases
In this section, we elaborate more on false positive (FP) cases observed through our privacy policy analysis. We take the three data types with the highest number of false positives as examples and discuss the potential reasons for the false positive results.
Account info
The account information in this study refers to certain sensitive personal data that constitutes an Android user’s account, such as the username and email address when the user sets up his/her mobile device for the first time. According to our analysis outcomes, account information ranks at the top by the number of false positive cases observed.
One possible reason for these FPs may be the too-broad scope of account information based on the knowledge of pre-trained LLMs. For example, many third-party apps or services, e.g., social networks and banking apps, request users to create their own accounts. SDKs connecting those apps and services may access domain-specific account details rather than the Android user information, which our adopted LLM usually fails to distinguish these two types of data. One statement in the privacy policy of the SDK “clevertap.android.sdk” could support this explanation, as it describes “personal info” to be collected as “The way users refer to themselves, such as their first or last name, or nickname” in its privacy policy. Our NLI model-based analysis determines that the Android account information is mentioned although the “first or last name” and “nickname” only refer to the user details of the in-house services.
Another reason leading to the high FP occurrences might be because the pre-trained LLMs tend to hold a preservative stance when considering account information and adopt a too broad sense of personal data into this scope. For example in our analysis of the privacy policy of SDK “com.razorpay.checkout”, our adopted model produces a high confidence score when determining whether the sentence “Device or other IDs, Device or other IDs” implies users’ account information. In that case, our adopted model likely considers the account information under the scope of “other IDs” mentioned in the policy text.
Network info
The network info in this study refers to the diverse data of the connected network configuration, which could be jointly used for user profiling purposes. However, this term may be too coarse and ambiguous for the LLMs. FPs in this data type are caused by our adopted model’s misunderstanding, which mistakenly considers any explicit usage of the network or Internet as an access to network information. For example, in an SDK named “com.zendesk.chat”, a statement about collecting users’ location “Location is not resolved on the device, but it is resolved on the backend based on the client’s IP address” is detected to collect network information, too.
SMS
The SMS in this paper refers to the data obtained through invoking SMS-related resources and functionalities, contributing to the third-highest number of FP cases. Our manual analysis finds that the adopted LLM manages to understand the exact literal meaning of SMS. However, we suspect that our model behaves too preservative to extend the scope of SMS to all sorts of messages. For example, during our analysis of the SDK “com.zendesk.messaging”, we find that the LLM misunderstands the difference between e-mail and SMS messages. It produces a positive prediction for collecting SMS although the privacy policy means to send e-mails to the users for customer service purposes.
A.4. Writing into System Settings and Pre-installed Apps
Writing into System Settings
According to the Android developer documentation, an app is requested to pose WRITE_SETTINGS permission to add or modify an entry into the system settings. That permission is under the protection level signature. For that reason, only apps signed with the platform signature can obtain that protection level and proceed to system settings modification. These apps also have to be system apps, e.g., installed in “/system/app/” or “/system/priv-app/” directory, to request the WRITE_SETTINGS permission.
However, our manual inspection of AOSP source code finds that the aforementioned permission is not the only way to enable an app to write or modify system settings. When an app attempts to write into system settings, the OS will check the protection level that the requesting app has attained. If any one of the following five tags is found from the app’s ApplicationInfo metadata, namely “signature” (with the tag value of 2), “appop” (40), “pre23” (80), “preinstalled” (400), and “role” (4000000), the OS will allow the requesting app to write into the system settings. Among the five flags, the preinstalled flag is set to all pre-installed apps by the OS.
Pre-installed Apps
According to Google’s documentation for system developers, there is no restriction posed to system developers in selecting pre-installed apps on a particular device. In other words, pre-installed apps do not have to apps with shared uid (i.e., user ID) and/or signed with the platform signature, such as android.uid.shell, android.uid.bluetooth, android.uid.phone, etc., but also be arbitrary third-party apps despite this practice is not encouraged by Google (google2024system, ).
According to the definition of protection level in AOSP source code (you may refer to the code “R.attr#ProtectionLevel” in the AOSP repository), “preinstalled” is a flag that is automatically granted to an arbitrary app pre-installed on the system image. An app with a “preinstalled” flag is given privilege to perform certain sensitive operations. Apart from writing into the system settings, there are some other privileged permissions can be automatically granted to pre-installed apps, including but not limited to AMBIENT_WALLPAPER, MANAGE_EXTERNAL_STORAGE, MANAGE_MEDIA,
READ_DEVICE_CONFIG, etc.
Note that pre-installed apps are not equivalent to “preloaded apps”. Preloaded apps can be installed in an ordinary location (e.g., /data/app/) or system directory (e.g., ‘‘/system/app/’’ and ‘‘/system/priv-app/’’), but pre-installed apps must be installed in a system directory. Generally, pre-installed apps stand for preloaded apps granted “preinstalled” privilege.
A.5. List of Tested Devices
No. Manufacturer Model
1 Xiaomi Xiaomi 11
2 Oppo Find X3
3 Vivo X70 Pro +
4 Honor Honor 50
5 OnePlus OnePlus 9
6 Samsung S21
7 Huawei P40
8 Lenovo Z6
9 Google Pixel Pixel 4 (US version)