### v2

# Assessing Privacy Compliance of Android Third-Party SDKs

Abstract
Third-party Software Development Kits (SDKs) are widely adopted in Android app development, to accelerate development pipelines and enhance app functionality effortlessly. However, this convenience raises substantial concerns about unauthorized access to users’ privacy-sensitive information, which could be further abused for illegitimate purposes like user tracking or monetization. Our study offers a targeted analysis of user privacy protection among Android third-party SDKs, filling a critical gap in the Android software supply chain. It focuses on two aspects of their privacy practices, including data exfiltration and behavior-policy compliance (or privacy compliance), utilizing techniques of taint analysis and large language models. It covers 158 widely-used SDKs from two key SDK release platforms, the official one and a large alternative one. From them, we identified 338 instances of privacy data exfiltration. On privacy compliance, our study reveals that more than 30% of the examined SDKs fail to provide a privacy policy to disclose their data handling practices. Among those that provide privacy policies, 37% of them over-collect user data, and 88% falsely claim access to sensitive data. We revisit the latest versions of the SDKs after 12 months. Our analysis demonstrates a persistent lack of improvement in these concerning trends. Based on our findings, we propose three actionable recommendations to mitigate the privacy leakage risks and enhance privacy protection for Android users. Our research not only serves as an urgent call for industry attention but also provides crucial insights for future regulatory interventions.
Index Terms:
mobile, privacy assessment, taint analysis.
I Introduction
The advent and ubiquity of third-party Software Development Kits (SDKs) in Android app development herald a profound shift in the mobile software development paradigm, transforming efficiency and accelerating innovation. These SDKs provide pre-built reusable code that offers a broad spectrum of functionalities, from fundamental features like user authentication [1] and data encryption [2] to specialized components like analytics [3], advertisement delivery [4], and user interface enhancements [5]. The open-source nature and global reach of the Android ecosystem further fuel the proliferation of these SDKs.
Integrating an SDK into an app is often as simple as declaring dependencies in the build configuration file, thus drastically reducing development time, and allowing developers to focus on app-specific features. Nowadays, the majority of Android apps incorporate multiple third-party SDKs, with an average app integrating approximately 8.6 SDKs [6, 7]. However, this widespread dependency also introduces significant security and privacy concerns. The complexity of the software supply chain and the opaque data handling practices of some SDKs can lead to privacy risks and security vulnerabilities, either due to malicious intent or unintentional oversights. These can be exploited by malevolent actors or lead to inadvertent privacy violations, as evidenced by recent high-profile incidents related to third-party SDKs [8, 9].
The privacy issues caused by SDKs could be further magnified in the pre-installed apps on Android devices. These apps are bundled by device manufacturers or network carriers, and therefore, often have an extremely great user population. They come with high privileges and are typically non-removable, such that an SDK embedded by them can stealthily perform higher-privileged operations that normal apps are unable to. An illustrative example is an app named “Mobile Services Manager” pre-installed on certain Android devices, which was found to download apps without user consent due to a malicious third-party SDK embedded [10, 11].
Assessing privacy practices of SDKs includes two key aspects, i.e., data exfiltration and behavior-policy compliance (or privacy compliance). The data exfiltration involves the behaviors of locating sensitive data in publicly accessible places, such as the publicly accessible directories in file systems and the Internet. The opacity of SDKs’ operations and their “black-box” nature hinders meaningful assessments, leaving both end-users and app developers in the dark and elevating risks. Dishonest or malicious SDKs may disguise as their embedding apps to extensively collect privacy-sensitive data, while the end-users are unable to distinguish which party requests their privacy data. Incidents like the X-Mode controversy [12], where user locations were tracked by over 100 apps through embedded SDKs, underscore this point and serve as a stark reminder of the challenges associated with preserving user’s privacy in third-party SDKs.
On the regulatory front, data protection regulations have been put in place globally, such as the General Data Protection Regulation (GDPR) [13] and the California Consumer Privacy Act (CCPA) [14], which impose a stringent requirement on data processors. In response to this, software developers are required to provide a privacy policy to explicitly disclose what personal information is collected and how these data are handled by their software. Dissecting whether the data handling behaviors of SDKs comply with the claims in their privacy policies, which we define as behavior-policy compliance (or privacy compliance), becomes an urgent task.
Our Work
In this work, we conduct a large-scale study on the privacy practices of existing third-party SDKs. Our study covers 158 prevalently-installed SDKs from two mainstream release platforms, including Google Play SDK Index [15] and CAICT SDK Index [16]. The former is the official source of Android SDKs as a part of Google Play infrastructure, which lists the most widely used commercial SDKs in the Android app ecosystem [17]. The latter is claimed to be the largest release platform of mobile SDKs in China, serving as a complementary source of SDKs due to the absence of Google’s service in the Chinese market. Based on the statistical data provided by the two platforms, over 100 collected SDKs have been embedded by at least 100 apps with over 10 million installations, demonstrating the broad impact of this study.
Our study comprises two main strategies, i.e., conducting a data exfiltration assessment through static program analysis, and assessing privacy compliance through interpreting privacy policies. Specifically, we conduct a static taint analysis for tracking the flow of privacy-sensitive data within the SDK. We explore the traces that read and share privacy-sensitive data in public spaces, and accordingly investigate whether these traces can lead to potential privacy-sensitive data exfiltration. For the privacy compliance assessment, we employ a large language model (LLM) to analyze the SDKs’ privacy policies and thereby identify what data is requested, cross-check the data claimed in privacy policies with the data access behaviors caught during the taint analysis, and determine if an excessive collection or over-claiming issue exists.
Our study reveals critical privacy shortcomings in Android third-party SDKs. Astonishingly, 346 traces are found to read privacy data and share them in publicly accessible places. Among these traces, 338 are confirmed to be subject to data leakage. Only 109 out of the 158 examined SDKs provide privacy policies, among which approximately 37% over-collect private data and over 88% falsely claim their data collection scopes, seriously violating global data protection norms. Our re-inspection after 12 months shows no meaningful improvement in these practices over time. We also uncover a novel system settings injection issue: some malicious SDKs exploit pre-installed apps to share privacy-sensitive data into Android OS’s public system settings, effectively creating a unique user-unresettable identifier (UUI) accessible to any app. This unprecedented finding highlights an urgent need for improvement in SDK development practices, permission management, and privacy violation detection. To this end, we further propose three key strategic recommendations as mitigation.
Contributions
Our work makes the following key contributions:
  • Detecting Potential Privacy Data Exfiltration on a Large Scale. We collect 158 SDKs from two mainstream repositories and leverage taint analysis to track sensitive data flow, thereby exploring their potential data exfiltration. We find 338 data exfiltration evidenced by sharing obtained privacy data into public spaces. Our work highlights the substantial privacy risks stemming from the widespread use of third-party SDKs in the Android ecosystem, unveiling complex and often opaque data-handling practices.
  • Revelation of Privacy Compliance Landscape at the SDK level. Our study uncovers a startling level of data over-collection among third-party SDKs, with an estimated 37% of SDKs engaging in this practice. Through our re-inspection, we depict the trends of SDKs’ practices in collecting and sharing privacy data, and underscore the urgency of tackling privacy risks in the Android ecosystem at a fine-grained SDK level.
  • Discovery of Privacy Risk caused by Pre-installed Privileges. We find a system settings injection issue in the pre-installed apps, in which a malicious SDK can excessively read and share privacy data into the system settings. This enables an arbitrary pre-installed app with the malicious SDK to create uniquely unresettable identifiers and share them, violating the relevant regulations about privacy protection. It is the first time such issues have been revealed and systematically investigated.
II Backgrounds and Problem Definitions
II-A GDPR and Privacy Compliance
The General Data Protection Regulation (GDPR) [13], a landmark legislation enacted by the European Union (EU) in 2018, has profoundly transformed the landscape of data privacy laws within the region. Under GDPR, entities processing personal data, defined broadly as any data enabling the direct or indirect identification of individuals, are subjected to stringent regulations that underscore the importance of “lawfulness, fairness, and transparency”. Furthermore, GDPR provides explicit requirements for obtaining valid consent, necessitating that it is freely given, specific, informed, and unambiguous.
Privacy compliance, in the context of GDPR, necessitates a thorough examination of privacy policies to ensure adherence to the GDPR’s comprehensive principles and stipulations. It entails a meticulous inspection of data processing activities, thereby ensuring data is acquired and utilized only for explicit, legitimate purposes. Furthermore, the principle of “data minimization” obliges organizations to limit their data processing activities to the minimum necessary, thereby reinforcing the importance of data accuracy and relevance. This process also extends to an examination of storage practices, necessitating the validation that data is retained only for as long as necessary and is safeguarded against unlawful processing, inadvertent loss, or damage. Privacy compliance, therefore, functions as a crucial mechanism to enforce accountability, reinforcing the necessity for organizations to demonstrate, unequivocally, their adherence to GDPR principles.
II-B Android SDKs and Data Collection Practices
Android third-party SDKs represent a crucial component within the broader ecosystem of Android app development. These software SDKs offer pre-written code to app developers, enabling the efficient integration of different features and functions within their apps, ranging from user interface elements to complex data analytics capabilities. Data collection practices associated with Android third-party SDKs have been a subject of scrutiny within academic and industry communities. It has been observed that some SDKs, while providing utility functions, also engage in extensive data collection activities without explicit acknowledgment or consent received from the end-users. Furthermore, there are instances where data is transmitted to remote servers, raising concerns about data security and potential misuse. In light of this, there is an increasing need of a large-scale privacy assessment of SDKs.
II-C Objectives and Problem Definitions
This work aims to (1) detect potential exfiltration of users’ privacy data existing in the third-party SDKs, (2) explore the privacy compliance of developers in the SDK release, and thereby, (3) unveil the landscape of privacy protection in the Android ecosystem at the SDK level. To this end, we propose a compliance model to systematically assess the privacy practices of Android SDKs regarding personal information collection. We let $\mathbb{D}$ represent a set of diverse data considered as user privacy, and let $\mathbb{S}$ be the set of publicly released third-party SDKs. Then we define what privacy data an SDK claims to collect and actually collects as follows.
Compliance Disclosure ($\mathcal{C}$)
We let $\mathcal{C}$ denote the compliance disclosure of an SDK $s\in\mathbb{S}$. Specifically, it defines the set of privacy data claimed to access in the privacy policy of the SDK $s$. It shall also be assumed not to leak any privacy data to the public.
Implementation-level Practice ($\mathcal{P}$)
The compliance practice of an SDK concerns two types of operations in its implementation, i.e., collecting privacy data, and sharing the collected privacy data. We let $\mathcal{R}_{s}$ denote the scope of privacy data that would be read/collected by an SDK ${s}$ based on its implementation, in which $\mathcal{R}_{s}\subseteq\mathbb{D}$. We also let $\mathcal{U}_{s}$ denote the collected privacy data involved in any uploading/sharing operations, given that $\mathcal{U}_{s}\subseteq\mathcal{R}_{s}$. Thus, we have the compliance practice of an SDK ${s}$ as a tuple of its collecting and sharing behaviors, written as $\mathcal{P}_{s}=(\mathcal{R}_{s},\mathcal{U}_{s})$.
Next, we formalize three types of compliance issues concerned in our investigation.
Type I: Privacy Leakage
[…]
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},d\in\mathcal{U}_{s}\Rightarrow\,\,\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$ (1)
Besides improper sharing, recent developments in privacy protection regulations also pose stricter requirements in even reading privacy data on personal devices. Developers are required to not only declare a list of privacy data to be collected but also to ensure the actual collection operations are always consistent with what has been declared. Next, we formalize two types of inconsistency issues that could happen in privacy data collection.
Type II: Excessive Collection
Excessive collection concerns the privacy practice of an SDK that attempts to read more types of privacy data than what it requests/claims in its privacy policy. Excessive collection infringes on users’ right to be informed as the SDK collects privacy data without users’ awareness and consent. We define this type of risk as follows.
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\in\mathcal{R}_{s})\land(d\notin\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$ (2)
We remark that Google has not set a formal definition of “data collection” at the SDK level as what it does for third-party apps [18]. Thus, we refer to the guideline of GDPR Article 13 [13, 19] and take a conservative strategy by considering data accessed by SDKs as collected when evaluating the excessive collection, given that the data accessed by an SDK can flow into the host apps, which can then consume or send the data out.
Type III: Over-claiming
We assert that an SDK has an over-claiming issue if it is found to claim access to more types of privacy data than it actually reads. Over-claiming issues, although they may not substantially lead to privacy data exfiltration, seriously violate the data minimization principle stipulated in Article 5(1)(c) of GDPR. It is defined as follows.
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\notin\mathcal{R}_{s})\land(d\in\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvDash_{\mathcal{P}}\mathcal{C}_{s}.$ (3)
II-D Threat Model
We consider that the exfiltration of users’ privacy data caused by malicious SDKs should follow a unified routine. To achieve such exfiltration, the malicious SDK (the attacker) gains capabilities as follows:
  • The malicious SDK is released in a public platform/market and thus, can be embedded in an arbitrary app and installed on an Android device. Without causing an advantage to the attacker, that device is not rooted and is properly maintained with the latest security update installed.
  • The malicious SDK accesses users’ privacy data through the interfaces of the Android OS with permissions granted to the embedding app, i.e., all its collection of users’ privacy data must be realized on behalf of the embedded app.
  • The malicious SDK saves or shares the collected privacy data in a public space, i.e., there exists no party (including the embedding app) colluding with the malicious SDK in sharing the collected privacy data, or saving it in private storage.
III Methodology
III-A Approach Overview
Figure 1: The workflow of potential data exfiltration exploration and compliance assessment at the SDK level
We propose a four-phase approach and briefly illustrate it in Figure 1. In Phase 1, we collect 158 Android SDKs, including their binary files, privacy policies, and metadata from two release platforms. After that, Phase 2 involves analyzing the privacy policies using state-of-the-art LLMs to identify the types of privacy data requested by the SDKs. In Phase 3, we perform static taint analysis on the SDKs to examine their data collection behaviors at the implementation level, aiming to identify traces that constitute a complete privacy data reading and sharing process. SDKs caught to read and share user privacy would be prone to privacy leakage (i.e., type I issues). Finally, in Phase 4, we assess privacy compliance by cross-checking the data requested in the privacy policies with the actual data collection behaviors, aiming to detect excessive collection (Type II) and over-claiming (Type III) issues.
III-B Scope of Privacy Data
Considering most privacy data stipulates an explicit permission for apps to access, we mainly resort to the list of permissions [20] from Android developer documentation to define the list of privacy data. The list is further complemented with privacy-sensitive items recognized from the privacy or security change section of each Android OS release report, for example, the clipboard and a group of user-unresettable identifiers (UUIs) are recognized in the release document of Android 10 [21] and added to the scope of this study. In addition, we also take existing literature in personal information collection [22, 23, 24, 25, 26, 27] and app analysis [28, 29, 30, 31] into consideration to complement the list of privacy data. We thus include the usage of diverse sensors (e.g., motion and light sensors) into the list because they have been widely studied in the research community. As a result, we recognize 41 types of privacy data and categorize them into five groups. These 41 data types constitute $\mathbb{D}$ defined in Section II-C and formulate the scope of our compliance assessment. The list of recognized privacy data is shown in Table I.
TABLE I: List of 41 pre-identified privacy data types and the distribution of the 741 source methods
Data type & class Count Data type & class Count
(C1) Chip, Cellular & Peripheral Wifi 38
Call phone 41 Wifi MACs 5
Carrier info 5 Subtotal (245)
Cellbroadcast 2 (C3) Location & Sensors
External storage 18 Camera 54
ICCID 15 Location 87
IMEI 5 Misc. sensors 140
IMSI 3 Subtotal (281)
MEID 2 (C4) Media & Software Specific
Network info 6 Android ID 2
Phone status 1 App list 7
SD card serial 4 Audio record 22
Serial 4 Google Ad ID 3
SIM info 7 Media location 1
SMS 23 OAID 9
Telephone number 12 Screen record 1
Subtotal (148) Subtotal (45)
(C2) Wireless Communication (C5) Personal Data
Bluetooth 117 Account info 6
Bluetooth call 6 Browser bookmarks 1
Bluetooth ID 2 Calendar 2
Bluetooth MAC 36 Clipboard data 2
IP address 2 Contact list 9
SIP service (VOIP) 21 Contact log 2
SSID/BSSID 17 Subtotal (22)
Ultra-wideband 1 Total 741
Includes coarse location that is determined by network, and fine location that is jointly determined by network and GPS.
Categorization
We then categorize the identified 41 types of privacy data by three criteria:
  • We primarily group those privacy data that share a similar functionality or purpose. This is to ease our privacy policy based compliance assessment, because many developers may tend to only declare the data to be collected roughly by functionality (e.g., device identifiers, location, etc.) rather than accessing specific data or certain APIs.
  • We then refine our categorization by grouping those data that can be programmatically accessed through APIs in the same or similar classes based on the Android developer’s documentation. This is to facilitate our next step’s taint analysis as we need to collect as many involved APIs as possible to be the source methods, which we will detail in Section III-E.
  • We also consider the API permissions of collecting those privacy data into our categorization. Due to the historic changes in Android privacy mechanism are mainly reflected in permission rules of calling sensitive APIs, grouping personal data that involve APIs mentioned in the same privacy change can facilitate our discussion and gain insights if any non-compliance is observed. A typical example is that the invocation permission of most APIs to read device identifiers including serial, IMEI, and ICCID, are escalated in Android 10. We accordingly group these aforementioned data into one category.
Based on the criterion listed above, we further categorize them into five classes, namely chip, cellular & peripheral (C1), wireless communication (C2), camera, location, and other sensors, (C3), media and software specific (C4), and miscellaneous personal data (C5).
Ontology for Semantic Relationships
Considering the request of personal data in privacy policies may not always written in a precise manner as defined in the developer’s documentation, and sometimes certain privacy data items have multiple names or colloquial phrases, we refer to relevant research [32, 33] and propose an ontology to capture the semantic relationship among the 41 types of privacy data, and consequently to guide our assessment.
First, we manually collect alternative names or colloquial phrases for each type of identified privacy data and treat them as synonyms ($\equiv$). For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. Thus, all the synonymous terms of a privacy data type will be considered in our assessment.
[…]
III-C Phase 1: Collecting Android SDKs
Our study starts by collecting as many Android SDKs as possible from mainstream sources. Since our approach involves cross-checking SDK privacy disclosures with actual data collection behaviors at the implementation level, we gather not only SDK binaries (typically in AAR and JAR formats) but also the privacy policies and relevant metadata provided by the developers, when available.
Sourcing SDKs from Multiple Repositories
We identify the Google Play SDK Index as the primary source for collecting Android SDKs. Additionally, considering the uniqueness of mainland China’s Android ecosystem where Google Play infrastructure is absent, we use CAICT as a supplementary source. As a result, we download the latest version of 158 distinct SDKs as of October 2022, among which 139 are from the Google Play SDK Index and 19 are released on the CAICT website. These 158 constitute our previously defined $\mathbb{S}$ (Section II-C) for the later assessment. We also revisit the latest versions of the same group of SDKs in October 2023 and conduct an additional assessment, which we will detail in Section IV-D.
We remark that, at the time of this study, both the Google Play SDK Index and CAICT have not enforced strict regulation or code auditing mechanisms like the Google Play App Store. As a result, there is no guarantee that the SDKs from these sources include comprehensive documentation or privacy policies required by regulations such as GDPR.
Crawling Meta Information and Privacy Policies
Next, we implement a crawler with Selenium [34] to retrieve meta information and privacy policy links from the two sources. The meta information includes the SDK name, provider (developer’s identity), Maven IDs, and functional categories. Such information will be jointly used to uniquely identify SDKs during our subsequent analysis. We then collect the privacy policies from the retrieved links. We will use them to extract the list of personal data that the SDK requests.
The retrieval of privacy policies is found more complex than collecting meta information due to the lack of a compulsory standard for drafting them. For SDKs from the Google Play SDK Index, we rely on the links provided in the “data safety section” section if available, and download the corresponding webpages. Similarly, for SDKs from the CAICT website, we download the webpages from the “SDK privacy policy” link on the SDK details page. We use Google Translate to process all non-English text on these webpages.
Due to the limited availability of privacy policies, we only collected policies for 52 SDKs directly from the two platforms. For those do not provide privacy policy URLs, we manually searched the Internet using keywords like “disclaimer”, “policy”, “disclosure”, and “guidance” to look for any available privacy policies drafted by the SDKs’ developers. As a result, we identify privacy policy webpages for another 57 SDKs. In total, we obtained valid privacy policies for 109 SDKs (69.0% of all SDKs). The text from these webpages will be extracted for further analysis.
III-D Phase 2: Identifying Data Requested in Privacy Policies
After collecting the 109 privacy policies, our next step aims to identify the data requested by each SDK from them. In contrast to traditional app privacy policies that are written in the user-oriented language, SDK privacy policies are more developer-oriented and may contain many domain-specific terminology (e.g., ICCID, BSSID) and technical synonyms (e.g., Google Ad ID, which also appears as Advertising ID, or GAID). This uniqueness makes it challenging for existing methods [35, 32, 33, 36, 37, 38] to extract domain-specific entities accurately. To tackle these challenges, we leverage an LLM-based Natural Language Inference (NLI) model to infer the data collection practice from the privacy policies.
Figure 2: The workflow of our privacy policy analysis to find out what an SDK claims to collect
Adoption of NLI Model
NLI has proven effective in understanding natural language across different domains [39]. In this work, we use the state-of-the-art LLM-based NLI model named roberta-NLI [^1] The full name of the model is roberta-large-snli_mnli_fever_anli _R1_R2_R3-nli. We shortly write as roberta-NLI to save space. to identify requested data in privacy policies. By parsing sentences from these policies as premises, we input them into the model alongside hypotheses based on 41 pre-identified data types (refer to Section III-B), allowing the model to assess the entailment relationship between them. We briefly illustrate our approach in Figure 2.
There are two key challenges arose during the hypothesis formulation. The first challenge is the missing of a canonical writing standard of SDK privacy policies. Privacy policies often use developer-oriented language with technical synonyms, requiring us to manually gather synonyms (e.g., the advertising ID and GAID are interchangeably used to refer to the Google Ad ID) to ensure hypothesis completeness. Besides that, we also tailored hypotheses by matching different verb phases (e.g., “collect”, “read”) to each data type and its synonyms to maximize the entailment score when the collection of the inferred data type is mentioned in the privacy policy.
The second challenge is handling the complex context of SDK privacy policies, as the NLI model is not domain-specifically trained. For example, we found common sentence formats in privacy policies, such as “According to GDPR, all collection of device identifiers must be declared in advance.” Conventional inference models might incorrectly conclude with high confidence that this implies the collection of device identifiers, but it is irrelevant to the actual behaviors of the SDK being analyzed. To address this issue, we created three tailored hypotheses for each combination of verbal phrase and object, namely the positive hypothesis (shown as PH in Figure 2, e.g., “It will collect …”), negative hypothesis (NH, e.g., “It does not collect …”), and irrelevant hypothesis (IR, e.g., “It does not mention whether to collect …”). As a result, we have composed a list of 507 hypotheses covering the 41 pre-identified data types.
Inference of Data Requested
Given three hypotheses have been defined for each combination of verbal phrase and object, we only consider a data type claimed to be collected by an SDK if there exists at least one combination of the associated object (e.g., location) and applicable verbal phrase (e.g., request) satisfying the logical condition as follows:
$(\text{Score}_{p}>T)\wedge(\text{Score}_{p}>\text{Score}_{n})\wedge(\text{Score}_{p}>\text{Score}_{i})$
where $\text{Score}_{p}$ indicates the entailment confidence score for the positive hypothesis. An entailment inference requires $\text{Score}_{p}$ to be greater than a threshold value, written as $T$. Simultaneously, $\text{Score}_{p}$ must be greater than the scores for negative and irrelevant hypotheses (i.e., $\text{Score}_{n}$ and $\text{Score}_{i}$), indicating that the premise does not only “mention” the collection of that data type but also confirm that collection behavior is “performed” by the SDK itself.
Finding the Optimal Threshold
Intuitively, to infer a specific type of data being collected, the entailment prediction score ($\text{Score}_{p}$) is supposed to be greater than 0.5 to ensure it is always higher than the prediction of the neutral and contradiction hypotheses. However, our preliminary experiments observe a large number of false positives by setting the threshold to 0.5. For that reason, we perform a small-scale analysis to tune the threshold value from 0.5 and above until we find an optimal value that minimizes the number of false positive predictions. Specifically, we assume that Google should provide high-quality privacy policies for their own SDKs, and therefore we collected five distinct privacy policy shared by 13 Google in-house SDKs. We manually analyzed declared privacy data types from the five privacy policies to set a ground truth. We then fine-tuned the threshold value and discovered that the NLI model could achieve perfect prediction with a threshold of 0.73. For that reason, we adopt 0.73 as the optimal threshold in our large-scale privacy policy analysis.
III-E Phase 3: Exploring Privacy Collection and Leakages
We leverage the static taint analysis techniques to explore privacy collection behaviors and potential exfiltration in SDKs. We define a valid privacy data exfiltration is constituted of a collection action that queries one or more privacy data from the OS, and a sharing action that shares the obtained privacy data, directly or indirectly, to the public spaces including the Internet, system settings, and the public storage. Thus, we consider that an SDK engages in privacy collection if its implementation contains a data flow originating from a privacy data collection action. This privacy collection leads to exfiltration if the data flow ends at a sensitive sharing action.
Configuring Taint Analysis
We aim to collect all relevant APIs that contribute to a collection action and include them into the source methods for taint analysis. To this end, We refer to an existing research [40] which proposes a technique named SuSi, which classifies over 900 API methods that can potentially leak users’ privacy in Android 4.3 and earlier versions. Considering there have been multiple version’s evolutions since Android 4.3 that the API rules may have experienced significant change, we first map the APIs identified in [40] with the 41 types of privacy data. We iterate the developers’ documents of those APIs and filter out those APIs that are no longer valid in Android Open Source Project (AOSP) implementation. We then manually visit the developer’s document and AOSP source code to collect as many APIs that are associated with the new data types as possible. As a result, we collect 741 source methods. [^2] We count the number of methods by unique method signatures in this paper. The distribution of 741 methods corresponding to the 41 types of privacy data can be found in Table I.
TABLE II: A brief statistics of recognized sink methods
Purpose Class name Count
Network Transfer java.net.HttpURLConnection 1
java.net.URL 2
java.net.URLConnection 2
org.android.spdy.SpdyRequest 12
org.android.spdy.SpdySession 3
org.apache.http.client.methods.HttpGet 8
org.apache.http.client.methods.HttpPost 8
org.apache.http.params.HttpParams 1
File Saving java.io.Writer 5
java.io.FileOutputStream 1
System Settings android.provider.Settings 3
Total 46
Next, we retrieve all methods that perform a sharing action from Android SDKs. Although there are not many new APIs introduced after Android 4.2 that can be treated as sink methods, we find most sink methods studied in earlier literature including SuSi [40], such as logging and SQLite databases, are no longer valid in the latest Android versions and therefore be excluded in this study [41, 42]. In the end, we identify 46 methods from 11 classes and then configure them as sink methods. We present a brief statistics of the sink methods by their containing classes in Table II.
The majority of sink methods are relevant to data transfer via network interfaces, reserving 37 out of 46 identified method signatures. Besides that, six method are recognized because they can share sensitive data via public storage. The remaining three sink methods are counted because they can be used for modifying the system settings, which is known as an undocumented channel to share data among different parties on the device [43]. We note that writing contents into the system settings requires privileged permission and we will discuss it in Section V.
IV Evaluation
Our evaluation aims to assess the performance of our proposed NLI model-based privacy policy analysis and explore the privacy data collection and potential leakages in the SDKs. After that, we analyze the privacy preservation landscape at the SDK level by cross-checking the data requested in the privacy policies and the data collected in SDKs’ implementation (refer to Phase 4 of Figure 1).
Implementation
We implement our privacy policy analysis in Python with PyTorch. The NLI model is provided by a Python package named Transformers. [^3] Available at https://huggingface.co/docs/transformers/index
Adoption of Automatic Taint Analysis
We learned that there are some off-the-shelf automatic static taint analysis tools for Android apps. Our preliminary investigation considers three open-sourced tools, namely FlowDroid [44], Mariana Trench [45], and AppShark [46]. We set up these three tools on our experimental PC, feed four best-selling non-game apps from the Google Play App Store, namely Facebook, Instagram, Snapchat, and TikTok, into these tools, and explore data flows simultaneously containing a privacy collection action and a sharing action. Our observation shows that AppShark performs the best in terms of time efficiency, success rate of analysis, and capability of handling inter-process communication (IPC). For those reasons, we adopt AppShark for the taint analysis of SDKs on a large scale. For each SDK, we implement a blank Android app to embed it and pass the app package to the automatic taint analysis program.
Evaluation Goals
Our evaluation aims to answer the following three research questions (RQs):
  • RQ1. Can our privacy policy analysis identify the privacy data claimed to be collected by SDKs on a large scale? What is the performance of our NLI model-based approach?
  • RQ2. Can our taint analysis find privacy data collection and sharing behaviors from the SDKs? How many tainted traces are detected? Can those tainted traces (really) leak users’ privacy?
  • RQ3. By cross-checking their privacy policies, what is the status quo of privacy data protection at the SDK level? Do the collected SDKs comply with their privacy policies?
IV-A RQ1: Privacy Policy Analysis Performance
TABLE III: Evaluation outcomes of privacy policy analysis (only data types requested are displayed)
* We use a broad sense of device identifiers to represent all hardware UUIs, including IMEI, MEID, IMSI, ICCID, and serial.
We perform NLI model-based analysis on the 109 SDK privacy policies and present the analysis outcomes in Table III. Our analysis identifies 214 data requests from privacy policies. Among the 214 data requests, IP address reserves the largest share, contributing to 47 (21.9%) occurrences. network information and location are the remaining data types among the top three. From the perspective of SDKs, the number of data types claimed by an SDK ranges from 0 to 6, with the average number 1.9.
We then conduct a manual inspection of our findings in the 109 privacy policies. As a result, we find 187 verified data requests covering 19 data types, and accordingly treat these requests as the ground truth in evaluating our analysis approach. The results demonstrate that our approach can effectively analyze privacy policies on a large scale, achieving a promising precision of 87.4% and an F1 score of 89.5%. (see Table III).
We find our NLI model-based analysis achieves perfect precision in detecting data collection for 10 data types despite 27 false positive cases being observed, i.e., wrongly recognized by the NLI model. Our evaluation outcomes show that account information (6), network information (5), and SMS (4) are the top three data types involved in false positives, indicating the model struggled to infer data collection for certain types. These errors likely stem from the model’s misinterpretation of terms and data type scopes. For example, although our NLI model correctly recognizes SMS, it fails to differentiate SMS and other messaging methods like emails, leading to high-confidence false positives while dealing with sentences like “we may send email to the users…” as SMS collection.
IV-B RQ2: Privacy Collection and Potential Leakages at the SDK Level
Our static taint analysis identified 2318 instances of privacy data access across 27 of the 41 pre-identified data types. Out of 158 SDKs, 95 (60.1%) were found to read privacy data, with 34 (21.5%) having at least one tainted trace. The analysis revealed that 346 (14.9%) of these accesses were tainted, suggesting potential privacy leakage into public spaces. We present the taint analysis outcomes by data types in Table IV.
TABLE IV: A brief statistics of the taint analysis results (only data types called and/or tainted are displayed)
Distribution of Data Collected & Shared
As shown in Table IV, location and sensors’ data (category C3) reserves the greatest portion of collection actions, contributing 1002 out of 2318 (43.2%) caught collection actions. The remaining privacy data categories ordered by their number of caught collection actions are data for wireless communication (C2, 22.8%), data for chip, cellular and peripheral (C1, 17.9%), media and software-specific data (C4, 13.4%), and personal data (C5, 2.8%), respectively.
From the privacy data sharing perspective, the access of wireless communication data (C2) constitutes the largest part of the tainted flow, as 172 out of 346 (49.7%) traces are tainted. The numbers of sharing actions involving data in categories C3 and C1 are following up, reserving 37.6% and 12.7% of all sharing actions, respectively. No sharing actions about data in categories C4 and C5 are observed. From the identifiability perspective, we find that most privacy data involved in sharing actions are referred to as personally identifiable information (PII) in relevant literature [25, 26, 47, 48].
We remark that whether SDKs can collect privacy data eventually depends on the permission and privilege of their embedding apps, and tainted traces do not automatically imply the exfiltration in runtime. A manual validation is performed to exclude false positive tainted traces. We will elaborate more on them in the remainder of this section.
1public static String a(Context context) {
2 ConnectivityManager cm = (ConnectivityManager) context.getSystemService(Context.CONNECTIVITY_SERVICE);
3 NetworkInfo info = cm.getActiveNetworkInfo();
4 if (info.isConnected()) { return "wifi"; }
5 ...
6 try {
7 return info == null? "" : info.getTypeName()+"-"+info.getSubtypeName()+"-"+info.getExtraInfo().toLowerCase();
8 }
9 catch(Exception e) { return ""; }
10}
Figure 3: A false positive example of SSID reading (line 7)
Potential Privacy Data Leakage (Type I Issues)
During our analysis, we noticed that static taint analysis cannot well determine the execution paths in conditional and branched code. If any tainted trace is found spanning over an unreachable, dead, or logically infeasible code, we exclude the tainted trace from the statistics of potential leakages. In Figure 3, we provide an example code reverse-engineered from one of our examined SDKs named “mipush”. We find it attempts to read the SSID of the connected WiFi but is later assessed by us as a false alarm. The SSID reading action (can be realized by the method getExtraInfo (line 7) until Android 10) would only take place if the device is not connecting to WiFi (after line 4), therefore we determine this is a logically infeasible code regarding privacy data access. In this study, we manually exclude the false alarms from the 346 traces and accordingly identify 338 (97.7%) logically valid traces, among which each trace indicates an occurrence of privacy leakage issue.
For the sharing operations, we observe that 332 out of 338 (98.2%) valid traces are tainted at network interfaces. The six remaining traces are tainted at system setting APIs. Compared with sharing privacy data through the Internet, we remark that writing privacy data in system settings is a more worrying issue because it exploits the privilege of pre-installed apps and is capable of sharing user privacy even without reading any privacy data from the Android OS. We will detail this in Section V.
IV-C RQ3: SDK-Level Privacy Compliance
We resort to cross-checking the privacy policies with the actual data collection behaviors of the collected SDKs to understand the status quo of privacy compliance at the SDK level. We remark our assessment only focuses on the 109 out of 158 (69.0%) SDKs that provide privacy policies, and the 27 data types that appear in the data collection behaviors observed from the taint analysis. We present our assessment results in Table V.
TABLE V: Assessment of privacy compliance by data types
Excessive Collection (Type II Issues)
Our evaluation finds that 40 out of 109 (36.7%) SDKs collected excessive privacy data compared with their privacy policies (e.g., an SDK collects personal identifiers but does not declare them in its privacy policy). Among the 40 SDKs that excessively collect users’ privacy, the number of undeclared data types ranges from 1 to 9. Each SDK has an average of three undeclared data types. As shown in Table V, 22 out of the 41 pre-identified data types (81.5%) have been excessively collected by the examined SDKs. The app list, clipboard data, and location are the top three most excessively collected data types. Specifically, 22 SDKs (20.2%) collect the app list without any declaration in their privacy policies, followed by clipboard data and location that are excessively collected by 11 (10.0%) SDKs. By jointly considering data collection behaviors (refer to Table IV), our findings show that the app list and clipboard data are the two most stealthily collected data types. Only 4 SDKs collecting the app list have mentioned it in their privacy policies (leading to 22 excessive collection behaviors). None of the 11 caught SDKs have declared to read the clipboard data.
Over-claiming (Type III Issues)
Apart from the excessive collection, we find that the over-claiming issue in privacy policies is far more pervasive in the tested SDKs. In our study, 96 out of 109 (88.1%) SDKs are found to claim more data types in privacy policies than what they (really) need. The IP address, network information, and MEID are the top three data types that have been over-claimed in the tested SDK privacy policies. Especially, 44 out of 109 (40.4%) SDKs that provide privacy policies over-claim their access to IP addresses. Such pervasive claims, again, reflect the worrisome landscape of privacy compliance at the SDK level.
Lessons Learned
Our study reveals two key findings. First, some SDK developers use vague or broad terms like “identifiers” or “device information” instead of specifying the exact data being collected, leading to differing interpretations. For instance, while most developers use “identifiers” to indicate device identifiers, some refer to advertisement identifiers, causing over-claims for device identifiers and under-claims for Android ID or Google Ad ID. Second, although many SDKs declare access to both GPS and IP addresses for location data, our taint analysis shows that most SDKs only use location APIs, leading to over-claiming cases of IP addresses. This lack of clarity can confuse users about what data is actually collected, reducing transparency in privacy policies.
IV-D Privacy Compliance Re-inspection
On the basis of our large-scale analysis performed on 158 SDKs collected in October 2022, we re-visited the two sources and collected the same group of SDKs in their latest release version as of October 2023 and conducted a cross-version analysis. Table VI depicts a summary of our re-inspection results.
TABLE VI: Changes in the tainted traces after 12 months (only tainted data types are displayed)
Data type & class Changes in tainted traces Data type & class Changes in tainted traces
[…]
$\dagger$ We use “new” to indicate data types observed re-inspection for the first time.
Trends of Privacy Collection Behaviors
We find 311 traces tainted from the latest version of 158 SDKs. Compared with our initial assessment, we observed 35 (10.1%) fewer traces that attempt to collect and share user privacy data in public spaces.
The change between the outcomes of the two assessments mainly falls in the sources of tainted traces. As shown in Table VI, we find obvious growth from the sharing of data relevant to media and installed software (C4), reserving over 1/3 of the total caught traces. Among the 105 tainted traces, reading and sharing the list of installed apps ranks the top among all sources, with 53 traces caught. Although the access to data in that category has been pervasively observed in our initial assessment, our re-inspection did not catch any traces of sharing such data.
Additionally, we find that the collection behaviors for location and SSID/BSSID have significantly declined, marking the two greatest decreases. There is also an overall reduction in the collection of UUIs, including data types in categories C1 and C2. Upon reviewing the historical changes in permission requirements, we note that most data types showing decreased collection behaviors had their access rules significantly updated in Android 10 [21], which imposed strict restrictions on third-party apps’ access to personally identifiable information. In contrast, we find the growth of tainted traces for the app list. Although the access to app list has also been restricted in a future release, these results still seriously concern us since the latest version of SDKs implies their persistent attempt to excessively and stealthily collect user privacy data.
V System Settings Injection Issues
Figure 4: Information flow demonstrates how the device ID is accessed and saved to the system settings by one of our analyzed SDKs (the obfuscated code is obtained from reverse engineering and has been simplified to save space)
TABLE VII: System setting entries found that contains a UUI
SDK Entry name Description of the value
[…]
Baidu LBS com.baidu.deviceid.v2 Package name and Android ID, MD5 $\ddagger$
[…]
$\dagger$ depending on the permissions granted to the embedded app.
$\ddagger$ inferred description based on the reading logic in the implementation.
During our analysis, we find some traces from an SDK named “Alicloud push” are sunk at writing a value into the system settings. Specifically, two entries of system settings are involved, as shown in Table VII. Writing into the system settings requires privileged permission so that ordinary third-party apps are unable to do so. However, we find that SDKs being embedded by pre-installed apps can write arbitrary values into the system settings even without the user’s awareness. We name this finding a system settings injection issue, because it can be manipulated by an SDK to create and share user-unresettable identifiers (UUIs).
Although leaking privacy data through covert channels has been studied in existing literature [49], the system settings injection issues still seriously concern us as the caught SDK subtly evades accessing highly sensitive UUIs (e.g., IMEI, as discussed in  [49]) but attempts to generate UUIs on their own. The caught SDK then takes advantage of its popularity and abuses the “preinstalled” privilege by counting on itself to be embedded in pre-installed apps or even privileged system apps. By exploiting this issue, the system settings become a covert channel to enable privacy-sensitive data sharing to apps without privileged permission.
Figure 4 illustrates an example of system setting injection from our findings. The source code is obtained through reverse engineering. In order to obtain a UUI, it attempts to invoke a privileged API (line 5, left), query the existence of pre-written system properties (lines 6 and 10-15, left), and read the Android ID (lines 7 and 16-24, left) sequentially, as shown in the figure. Due to the first approach requests the embedded app to be a system app with “READ_PRIVILEGED_PHONE_STATE” permission, and the second approach is only viable when there exists a system app on the same device to write the two system properties in advance, we consider both them infeasible by default on nowadays Android devices. However, the third approach, i.e., reading Android ID through a system setting table named SECURE, does not request special permission. As a result, the SDK can at least obtain the Android ID to uniquely identify the device. It then performs a hash function over the obtained Android ID (lines 19-20, right) and writes the hash value into the system settings with a customized key (lines 25-27, right). This process, although it complies with the Android permission mechanism, transforms the value of resettable Android ID into a de facto UUI that can be accessed by any app that knows the customized key in the system settings.
Considering the exploitability of the pre-installed privileges, we carry out another round of investigation of all 158 SDKs to find if there exists any other potential privacy infringement by taking advantage of the system settings mechanism. As a result, we managed to find another two entries that have been read in an SDK named “Baidu LBS” (see Table VII). As their names imply, these two entries are obviously written by third-party privileged apps. We suspect the formation of the two entries should follow a similar routine, and accordingly reverse engineer the involved SDK and present our inferred descriptions of the two entries in Table VII. Unfortunately, we did not find any app to write these entries into the system settings. Who exploited this system settings issue remains unknown to us.
Impact
We notice that the two studied SDKs are widely adopted in third-party apps, especially in the Chinese market. Therefore, we conduct a small-scale study to explore the impact of the system settings injection issues. We collect devices from 9 OEM manufacturers (8 OEM devices purchased in the Chinese market [^4] These devices are manufactured by Honor, Huawei, Lenovo, OnePlus, Oppo, Samsung, Vivo, and Xiaomi. and 1 Google Pixel device purchased in the US market), download the packages of pre-install apps, and investigate whether the two aforementioned SDKs have been once embedded in the pre-installed apps. We observe a worrisome fact that except for the Google Pixel devices, all the tested devices have at least one pre-installed app using the investigated SDKs. According to the latest market report [50], these 8 OEM manufacturers reserve over 66% global Android smartphone market share. They also shipped over 94% new Android devices in China in 2023 [51]. This implies that almost all Android smartphones sold in the Chinese market may already pose a privacy leakage risk based on an assumption that the choice of pre-installed apps should remain the same across different models produced by a manufacturer.
VI Discussion
VI-A Causes of Privacy Mishandling
Privilege Free-riding
An SDK’s access to privacy data depends on the permissions granted to the apps it’s embedded in. However, third-party apps pre-installed by OEMs can obtain privileged permissions. A malicious SDK in such an app can exploit system settings to deploy privilege free-riding attacks, creating and sharing UUIs without users’ awareness. As reading system settings require no permissions, privilege free-riding significantly undermines Google’s efforts in restricting third-party app access to UUIs since Android 10.
Insufficient Regulations
It is well known that the release of Android apps is subject to strict regulations, with non-compliance in privacy preservation potentially leading to removal from app stores or fines. However, SDK releases lack similar fine-grained regulations and sufficient legislation to ensure compliance with personal data protection laws. Ironically, even some Google in-house SDKs do not have privacy policies listed on the SDK Index [^5] An example can be found at https://play.google.com/sdks/details/com-google-android-gms-play-services-auth. Currently, both repositories function merely as search engines for available SDKs, without a platform-level compliance checking mechanism to ensure complete permission disclosure.
VI-B Our Recommendations
Precautions of Privilege Abuse
We regard that the pre-installed privileges open an attack surface for a malicious SDK, specifically when the SDK uses app-differential Android ID to identify a device user, transforms it into a de facto UUI, and provides free access with all apps installed on the victim device. Based on this finding, we advocate Google restricting pre-installed apps from writing data into the system settings in the future. We also advise OEM manufacturers to minimize the number of pre-installed apps on their devices.
Strengthened Platform Regulation
Currently, both release platforms selected in this study are merely index websites, which do not host the SDK binaries and lack substantial enforcement. Developers voluntarily register SDKs on these platforms. For that reason, we call for strengthened regulation over the usage of Android SDKs. We also advocate platforms performing static analysis over the SDK binaries to exhaustively list requested permissions, especially for those at the dangerous level and above. This measure could address the black-box concern for app developers.
Transparent App Development
We advocate transparent management over the permission requested by the embedding SDKs. Thus, app developers can have their own discretion in granting permissions without compromising functionality. We are delighted to see Google’s efforts in improving the SDK’s governance. A new mechanism called SDK Runtime has been introduced in Android 13 that provides a dedicated execution environment for SDKs [52]. However, this mechanism is not compulsory and still faces many usability and functionality limitations at the moment of this work being performed. Besides that, how to ensure the cooperation of SDK developers remains a challenging issue, given many SDKs are released and maintained in non-Google platforms.
VI-C Limitations
Absence of Dynamic Testing
While our static taint analysis approach is effective, dynamic testing can improve efficiency by automatically identifying and excluding logically invalid tainted traces, reducing manual validation efforts. However, dynamic SDK testing is challenging due to inadequate documentation, undisclosed interfaces, and highly event-driven/credential-required logic. For example, some finance and customer service SDKs require a valid license or token to function. Satisfying all functional prerequisites of the collected SDKs can be another challenging task and therefore is out of the scope of this study. We plan to explore potential dynamic testing solutions in the future.
Coverage of Targeted Methods
Our analysis covered 741 API source methods and 46 sinks. Although we frequently update the configuration during this study, we may still miss some data exfiltration channels as new APIs can be introduced along with OS evolution and emerging hardware.
Elusive and Ambiguous Disclosure
Although developers should clearly disclose personal data collection, we find that elusive or ambiguous phrasing widely exists in many SDK’s privacy policies, leading to potential false negatives in our NLI model-based analysis. We aim to explore the latest advancements in LLMs and improve our approach to minimize such impacts.
VII Ethics Considerations
We have reported all our findings to the relevant parties and we also have kept them confidential for at least 90 days. For each SDK that found a compliance issue, we manually searched the contact left by its developers and e-mailed our findings since our first round of SDK investigation. Although we observed slight mitigation from re-inspection, we did not receive positive responses regarding non-compliance concerns from the developers that we have attempted to contact. We also expressed our concern to Google about the system settings injection issues (to be detailed in Section V) and suggested removing the pre-installed privileges in writing system settings in the future release. Google replied and suggested that it is at the device manufacturer’s discretion to select and grant pre-installed privileges to third-party apps.
VIII Related Works
Privacy Compliance Evaluation
Numerous works focus on detecting inconsistencies between the privacy practices of an application and its descriptions or its privacy policies [53, 54, 55, 56, 57, 58, 59, 60]. This line of research begins with automatic risk assessment of Android apps, such as WHYPER [56] and AutoCog [57] that leverage NLP techniques to evaluate the necessity of required permissions based on apps’ descriptions. Gorla et al. [54] study mismatches between app descriptions and app behaviors by clustering Android apps based on their description topics and identifying outliers in each cluster with respect to their API usage. Recent work extends the scope of privacy compliance evaluation to examining third-party data harvesting, user consent, service misconfigurations, and quality of privacy policies at the app level. XFinder [27] uses dynamic analysis and NLP techniques to parse terms-of-service and extract data-sharing policies. Nguyen et al. [61] study user consent verification through traffic analysis and ablation experiments. Zhang et al. [62] highlight the issue of analytical service misconfigurations, while Yu et al. [63] scrutinize the apps’ privacy policies with regard to the inclusion of third-party libraries. Andow et al. [32] propose POLICHECK that identifies ambiguous or omitted disclosures of third parties. Pan et al. [64] study the automated generation of privacy policies for mobile apps and unveil that defective compliance widely exists in the automatically generated privacy policies. Zhao et al. [65] study privacy compliance of third-party libraries in Android through the lens of pattern matching in privacy disclosure. Compared with it, our work investigates SDKs, which are more complex in implementation and business logic. Our study also resorts to a language model in privacy disclosure analysis that does not request manual data annotation and model training. Parallel research by Xiao et al. [66] investigate compliance of third-party libraries by assessing the consistency between API documents and dynamic analysis results. In comparison, our work focuses on compliance of SDKs that contains more complicated implementation that hinders from effective dynamic analysis, and meanwhile our study examine the privacy disclosure by directly interpreting the privacy policies, which are written in natural language without a standard format.
Privacy Leakage Assessment
There is a substantial body of literature addressing various aspects of privacy concerns in the Android ecosystem [28, 67, 68]. Meng et al. [43] study user-unresettable identifier safeguards on a wide range of Android devices. He et al. [69] utilize dynamic analysis to explore the leakage of permission-related data from third-party libraries in Android apps Ekambaranathan et al. [70] investigate data usage and disclosure in children’s app. Liu et al. [71] examine data leakage from nine analytic libraries across 300 apps, using both static and dynamic analyses. Razaghpanah et al. [72] detect third-party advertising and tracking services via dynamic analysis of network traffic data.
IX Conclusion
In this study, we investigate the intricate privacy challenges associated with the widespread use of third-party SDKs in Android apps. Our large-scale analysis involved an evaluation of 158 third-party SDKs, prevalent across Android applications, with an approach intertwining rigorous taint analysis for data collection practices and LLM-aided interpretation of privacy policies. The findings were startling, with 338 potential privacy leakages detected from 158 SDKs. Besides that, our study reveals that less than 70% of examined SDKs provide privacy policies, among which approximately 37% SDKs are found indulging in the over-collection of privacy data, signaling a clear violation of privacy norms. Our findings suggest the need for stricter regulations, improved development guidelines, and more transparent permission management to reduce privacy risks in Android apps.